### 12 Mar 2024 v1

Partial Identification of Individual-Level Parameters Using Aggregate Data in a Nonparametric Binary Outcome Modelthanks: I am especially grateful to Edward Vytlacil and Isaiah Andrews for their guidance and advice. I thank Alberto Abadie, Anna Mikusheva, Hyungsik Roger Moon, Elie Tamer, and participants in the MIT econometrics lunch seminar for their helpful comments. All errors are my own.
Sarah Moon Department of Economics, Massachusetts Institute of Technology, sarahmn@mit.edu
1 Introduction
Researchers frequently use publicly available data in policy analysis, which is usually provided at the aggregate level due to individual privacy concerns. For example, statewide standardized exam results are reported in the form of school-wide or school district-wide pass rates, as opposed to pass result and demographic information (e.g. gender, race, family income) for each individual student. Election data is available as vote shares and voter demographics over voting districts, as opposed to the vote and demographics information for each individual voter. However, it is well known that aggregate variables can be related in ways that are different from the same variables at the individual level, [^1] Firebaugh (1978) notes that “aggregate” and “individual” are relative terms, as individual variables correspond to the individual unit of analysis and need not correspond to an individual person. In the econometrics literature, the terms “macro” and “micro” are also often used. a problem known as an ecological fallacy. [^2] This problem has also been referred to as the problem of aggregation (Stoker, 1984).
In this paper I consider the problem of identifying linear combinations of conditional mean outcomes, $\mathbb{E}[Y_{i}|X_{1i},\dots,X_{Li}]$, when individual-level outcome $Y_{i}$ is binary and the only data that is observed is the marginal distribution of each individual-level variable over many groups, which I call aggregate data. I develop a partial-identification methodology that constructs sharp bounds by solving an optimization problem that considers all underlying joint distributions of individual-level covariates that are consistent with the observed marginal distributions. I also derive closed-form bounds on each conditional mean outcome $\mathbb{E}[Y_{i}|X_{1i},\dots,X_{Li}]$. Finally, I show how further restrictions on the underlying data generating process, like shape restrictions, can be incorporated into the optimization problem to obtain sharp bounds under those restrictions. Since I do not observe individual-level joint distributions, I develop valid inference procedures on the identified set using marginal information only.
To demonstrate how informative these bounds can be, I apply this methodology to several different simulated aggregate data sets, calibrated to a Rhode Island standardized exam score data set used later in an empirical application. I find that bounds become narrower on specific conditional outcomes when the observed , and wider when the observed . Imposing monotonicity shape restrictions in the empirical application helps narrow bounds on test score gaps, but bounds remain wide. This suggests that obtaining useful individual-level results from aggregate data alone requires further, stronger assumptions that should be carefully justified by the researcher.
The ecological fallacy has been studied across many fields over the past several decades, beginning with the seminal work of Robinson (1950), Duncan and Davis (1953), Theil (1954), and Shiveley (1974). These early methods demonstrated issues with interpreting parameters from regressions with aggregate data as individual-level effects, and provided simple bounds on individual-level effects with a binary outcome and single binary covariate. Since then, research has been conducted on consideration of the identification of parameters (Stoker, 1984, 1986), testing for aggregation bias (Lee et al., 1990), empirical illustrations of inconsistent results when using aggregate as opposed to individual-level data (Hsiao et al., 2005), and many other topics in the sciences and social sciences beyond economics.
While the problem of aggregation has been widely acknowledged as a major issue by the literature, to the best of my knowledge much of the literature is concerned with point identification of individual-level parameters (King et al., 1999; Rosen et al., 2001). [^3] One exception is Jiang et al. (2020), who develop partial identification techniques in a setting with two binary covariates and linear individual-level relationships. When only aggregate information is available, knowledge about parameters at the individual level is limited, and point identification is often not achieved without further assumptions. Imposing more assumptions may allow for precise results, but such assumptions may be less plausible. For example, one popular method for analyzing aggregate data is the ecological inference method (King, 1997), used often in political science studies of elections (Burden and Kimball, 1998). This method relies on many assumptions, like imposing individual-level joint distributions and a lack of bias introduced by the aggregation, that often fail to hold in applied settings (Tam Cho, 1998; Cho and Gaines, 2004).
Even without these strong assumptions, partial identification is still possible and may reveal useful insights about the size or range of magnitude of parameters of interest. Partial identification in various contexts has been widely studied in the econometrics literature, especially over the last thirty years. [^4] See Tamer (2010), Ho and Rosen (2015), Molinari (2020), and Kline and Tamer (2023) for detailed surveys of partial identification in economics. Relevant to this paper is partial identification with data combination, since when combining two different data sets data the joint distribution between the data are unobserved (Cross and Manski, 2002; Molinari and Peski, 2006; Ridder and Moffitt, 2007; Fan et al., 2014, 2016). Relative to this literature, I assume that the joint distribution between every variable is unobserved and provide valid inference methods.
The rest of the paper proceeds as follows. Section 2 presents identified sets on the parameters of interest under a few different assumptions. Section 3 develops consistent estimation and inference procedures. Section 4 presents and discusses results from an empirical application using standardized exam data and simulation exercises calibrated to the dataset. Section 5 concludes.
2 Identification
2.1 Identified set
Let $(Y_{i},X_{1i},\dots,X_{Li},G_{i}),i=1,\dots,n$ be an i.i.d. sequence of random variables. [^5] I assume individual-level variables are i.i.d. for the sake of simplicity; this assumption can be relaxed and the consistency and inference results of Section 3 will still hold. Suppose outcome $Y_{i}$ is binary, with $Y_{i}\in\{0,1\}$ and suppose $G_{i}\in\{1,\dots,G\}$ denotes the group of individual $i$. Further assume covariates $X_{i}\equiv(X_{1i},\dots,X_{Li})^{\prime}$ is discrete with known finite support $\{x_{k}\}_{k=1}^{K}\subseteq\mathbb{R}^{L}$. [^6] The case with multinomial or continuous outcome and continuous covariates is beyond the scope of this paper.
The goal is to construct bounds on linear combinations of $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$, $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$, for given weights $\{\lambda_{k}\}_{k=1}^{K}$. For example, if we are interested in the average marginal effect of changing $X_{i}$ from $x_{k_{1}}$ to $x_{k_{2}}$ on $Y_{i}$, we would choose $\lambda_{k_{2}}=1,\lambda_{k_{1}}=-1,$ and $\lambda_{k}=0$ for all other $k$. I will construct identified sets using only expressions for $\mathbb{E}[Y_{i}|G_{i}=g],\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g],$ and $\mathbb{P}[G_{i}=g]$ for every $\ell=1,\dots,L,k=1,\dots,K,g=1,\dots,G$; the sample equivalents of these parameters are observed in aggregate data.
For example, in a data set of standardized exam results and demographics, $G_{i}$ denotes student $i$’s school district, $Y_{i}$ is an indicator for whether student $i$ passed the exam or not, and $X_{i}$ are student $i$’s demographics. We observe (sample estimates of) the pass rate for every school district $\mathbb{E}[Y_{i}|G_{i}]$, demographics of every school district $\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$, and the number of students enrolled in each school district, with which we can obtain (sample estimates of) $\mathbb{P}[G_{i}=g]$.
I summarize the previous statements in the following assumption:
Assumption 1.
For random variables $Y_{i},G_{i}$, and $L$-dimensional random vector $X_{i}$, suppose
  1. 1.
    $(Y_{i},G_{i},X_{i})$ are i.i.d.
  2. 2.
    $Y_{i}$ is binary.
  3. 3.
    $G_{i}$ is discrete with finite support $\{1,\dots,G\}$.
  4. 4.
    $X_{i}$ is discrete with finite support $\{x_{k}\}_{k=1}^{K}\subseteq\mathbb{R}^{L}$.
  5. 5.
    $(Y_{i},G_{i},X_{i})$ are latent; instead, we observe (sample analogs of) $\mathbb{E}[Y_{i}|G_{i}=g],\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g],$ and $\mathbb{P}[G_{i}=g]$ for every $\ell=1,\dots,L,k=1,\dots,K,g=1,\dots,G$.
Because $Y_{i}$ is binary, the law of total probability gives us
$\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ $=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\mathbb{E}[Y_{i}|X_{i}=x_{k}|G_{i}=g],$
$\mathbb{E}[Y_{i}|G_{i}=g]$ $=\sum_{k=1}^{K}\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]\mathbb{P}[X_{i}=x_{k}|G_{i}=g],$
$\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ $=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}\mathbb{P}[X_{i}=x_{j}|G_{i}=g].$
As in Cross and Manski (2002), these are the only relationships we can use to relate joint information of interest to the observed marginal information in the data without any further assumptions.
Let $\delta_{k}\equiv\mathbb{E}[Y_{i}|X_{i}=x_{k}]$, $\gamma_{kg}\equiv\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]$, and $\pi_{kg}\equiv\mathbb{P}[X_{i}=x_{k}|G_{i}=g]$ denote the unobserved parameters in the above equations. Then we can rewrite the equations as
$\delta_{k}$ $=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\gamma_{kg},$ (1)
$\mathbb{E}[Y_{i}|G_{i}=g]$ $=\sum_{k=1}^{K}\gamma_{kg}\pi_{kg}$ (2)
$\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ $=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}\pi_{jg}.$ (3)
Note that $\delta_{k},\gamma_{kg},\pi_{kg}\in[0,1]$ for all $k,g$. Combining this with equations (1), (2), and (3) above, we can define the identified set for $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$:
Lemma 1.
Suppose Assumption 1 holds. Given $\{\lambda_{k}\}_{k=1}^{K}\in\mathbb{R}^{K}$, the sharp identified set for $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ is given by
$D=\Bigg{\{}\sum_{k=1}^{K}\lambda_{k}d_{k}~{}\bigg{|}~{}0\leq d_{k}\leq 1~{}\forall k,\text{ and }\exists\,(p_{1g},\dots,p_{Kg})\in[0,1]^{K},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }d_{k}=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}~{}\forall k,~{}\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g,\\ \mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}~{}\forall\ell,k,g,\text{ and }\sum_{k=1}^{K}p_{kg}=1~{}\forall g\Bigg{\}}.$ (4)
Proposition 1.
$D=\left[\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]L_{g},\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]U_{g}\right]$, where
$L_{g}$ $\equiv\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},$
$U_{g}$ $\equiv\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},\text{ and}$
[…]
I defer all proofs to Appendix A. Proposition 1 states that we can equivalently express the identified set $D$ given by (4) as the weighted sum of solutions to bilevel optimization problems. This formulation is helpful because it suggests how computation of the lower and upper bound might be performed. In particular, solving for
$\min_{\{c_{kg}\}\in[0,1]^{K}}\underset{\{c_{kg}\}\in[0,1]^{K}}{(\max)}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ (5)
given a particular $\{p_{kg}\}\in P_{g}$ is a linear program. Thus we can solve for $L_{g}$ and $U_{g}$ by searching for the minimum and maximum respectively of the linear program (5) over all $\{p_{kg}\}\in P_{g}$. Not only is (5) a linear program, the following corollary shows (5) has a closed-form solution given any $\{p_{kg}\}\in P_{g}$:
Corollary 1.
For given weights $\lambda_{1},\dots,\lambda_{K}$ and any fixed $\{p_{kg}\}\in P_{g}$, relabel the indices $k=1,\dots,K$ so that $\frac{\lambda_{1}}{p_{1g}}\geq\dots\geq\frac{\lambda_{K}}{p_{Kg}}$, where if $p_{kg}=0$ we define $\frac{\lambda_{k}}{p_{gk}}\equiv+\infty$. Then
$\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$
is attained by letting
$c_{1g}$ =$=\begin{cases}\max\left\{0,\frac{\mathbb{E}[Y_{i}|G_{i}=g]-1+p_{1g}}{p_{1g}}\right\}&p_{1g}>0\\ 0&p_{1g}=0\end{cases},$
$c_{kg}$ =$=\begin{cases}0&\sum_{j=1}^{k}p_{jg}\leq 1-\mathbb{E}[Y_{i}|G_{i}=g]\\ \frac{\mathbb{E}[Y_{i}|G_{i}=g]-1+\sum_{j=1}^{k}p_{jg}}{p_{kg}}&\sum_{j=1}^{k-1}p_{jg}\leq 1-\mathbb{E}[Y_{i}|G_{i}=g]<\sum_{j=1}^{k}p_{jg}\\ 1&\sum_{j=1}^{k-1}p_{jg}>1-\mathbb{E}[Y_{i}|G_{i}=g]\end{cases},$
and
$\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$
is attained by letting
$c_{1g}$ =$=\begin{cases}\min\left\{1,\frac{\mathbb{E}[Y_{i}|G_{i}=g]}{p_{1g}}\right\}&p_{1g}>0\\ 1&p_{1g}=0\end{cases}$
$c_{kg}$ =$=\begin{cases}1&\sum_{j=1}^{k}p_{jg}\leq\mathbb{E}[Y_{i}|G_{i}=g]\\ \frac{\mathbb{E}[Y_{i}|G_{i}=g]-\sum_{j=1}^{k-1}p_{jg}}{p_{kg}}&\sum_{j=1}^{k-1}p_{jg}\leq\mathbb{E}[Y_{i}|G_{i}=g]<\sum_{j=1}^{k}p_{jg}\\ 0&\sum_{j=1}^{k-1}p_{jg}>\mathbb{E}[Y_{i}|G_{i}=g]\end{cases}.$
Although the solution to (5) has a closed-form solution, the optimization problem defined by searching for the minimum (or maximum) of (5) over all $\{p_{kg}\}\in P_{g}$ has a nonconvex objective. I suggest using existing derivative-free nonconvex solvers with a coarse grid of starting points to solve the optimization problem.
Remark 1.
Without further restrictions on the underlying data generating process, it will always be the case that $\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|G_{i}=g]\in D$. To see why, note that, inspecting the optimization problems of Proposition 1, for any $g$ and any $\{p_{kg}\}\in P_{g}$, letting $c_{kg}=\mathbb{E}[Y_{i}|G_{i}=g]$ for all $k=1,\dots,K$ satisfies the constraint that $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ because $\sum_{k=1}^{K}p_{kg}=1$. This is relevant for average marginal effect parameters because the weights $\{\lambda_{k}\}$ are such that $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|G_{i}=g]=0$ for each $g$, and we thus cannot rule out a zero average marginal effect.
2.2 Fréchet inequalities
In Section 2.1 I derived sharp bounds for a linear combination of $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$, defined using the solutions to optimization programs. I now show that if we are only interested in obtaining sharp bounds on each $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ parameter, we can obtain closed-form bounds using Fréchet inequalities.
The Fréchet inequalities, explicitly derived by Fréchet (1951), state that if there are $N$ events $A_{1},\dots,A_{N}$, it holds that
$\max\left\{1-N+\sum_{i=1}^{N}Pr[A_{i}],0\right\}\leq Pr\left[\bigcap_{i=1}^{N}A_{i}\right]\leq\min\left\{Pr[A_{1}],\dots,Pr[A_{N}]\right\}.$ (6)
Therefore, for $y$ in the support of $Y_{i}$ and $x_{k}$ in the support of $X_{i}$, it follows from the Fréchet inequalities that
$L_{g}(y,x_{k})$ $\equiv\max\left\{\mathbb{P}[Y_{i}=y|G_{i}=g]+\sum_{\ell=1}^{L}\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]-L,0\right\}$
$\leq\mathbb{P}\left[Y_{i}=y,X_{i}=x_{k}|G_{i}=g\right]$
$\leq\min\left\{\mathbb{P}[Y_{i}=y|G_{i}=g],\mathbb{P}[X_{1i}=x_{k,1}|G_{i}=g],\dots,\mathbb{P}[X_{Li}=x_{k,L}|G_{i}=g]\right\}$
$\equiv U_{g}(y,x_{k}).$ (7)
Note that we only require knowledge of the marginal probabilities of each random variable in order to be able to calculate the lower and upper bounds; joint probabilities are not needed. If we are only given $Pr[Y_{i}=y|G_{i}=g],Pr[X_{1i}=x_{1}|G_{i}=g],\dots,Pr[X_{Li}=x_{L}|G_{i}=g]$ and nothing else, then it is well known that the Fréchet inequalities are sharp; that is, they are the tightest possible bounds given the assumptions. Situations in which the Fréchet inequalities are not sharp include those in which we know variables are independent or if the (known) support of the variables is such that knowledge of a marginal probability provides knowledge of the joint probability.
For a given $k$ we can define bounds on $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$:
Proposition 2.
Suppose Assumption 1 holds. Let
$D^{F}\equiv\left[\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]L^{F}_{g},\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]U^{F}_{g}\right],$
where
$L^{F}_{g}$ $\equiv\sum_{k=1}^{K}\lambda_{k}\frac{L_{g}(1,x_{k})}{L_{g}(1,x_{k})+U_{g}(0,x_{k})},$
$U^{F}_{g}$ $\equiv\sum_{k=1}^{K}\lambda_{k}\frac{U_{g}(1,x_{k})}{L_{g}(0,x_{k})+U_{g}(1,x_{k})}.$
Then $D\subseteq D^{F}$, and if the Fréchet inequalities on $\mathbb{P}[Y_{i}=y,X_{i}=x_{k}|G_{i}=g]$ are sharp for all $y\in\{0,1\},k=1,\dots,K,g=1,\dots,G$ then $D=D^{F}$.
2.3 Additional shape restrictions
There may be situations in which the researcher is willing to assume polyhedral shape restrictions on the conditional expectation function over groups $\mathbb{E}[Y_{i}|X_{i},G_{i}]$. Examples of such shape restrictions include convexity, concavity, and monotonicity. [^7] See the reviews written by Matzkin (1994) and Chetverikov et al. (2018) for other examples of shape restrictions that have been used in econometric models.
I impose the shape constraints as an additional assumption:
Assumption 2.
For each $g=1,\dots,G$,
$S_{g}Y_{X,g}\leq 0,$
where $Y_{X,g}\equiv\left(\mathbb{E}[Y_{i}|X_{i}=x_{1},G_{i}=g],\dots,\mathbb{E}[Y_{i}|X_{i}=x_{K},G_{i}=g]\right)^{\prime}$ and $S_{g}\in\mathbb{R}^{s_{g}\times K}$ are fixed matrices.
We can simply add this shape restriction to the constraints of the optimization problems solving $L_{g}$ and $U_{g}$ in Proposition 1, as in Freyberger and Horowitz (2015), to obtain sharp bounds on $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$:
Proposition 3.
Suppose Assumptions 1 and 2 hold. Given $\{\lambda_{k}\}_{k=1}^{K}\in\mathbb{R}^{K}$ and $S_{g}\in\mathbb{R}^{s_{g}\times K}$ for each $g$, the sharp identified set for $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ is given by
$D=\left[\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]L_{g},\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]U_{g}\right],$
where, defining $c_{g}\equiv(c_{1g},\dots,c_{Kg})^{\prime}$,
$L_{g}$ $\equiv\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }S_{g}c_{g}\leq 0\text{ and }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},$
$U_{g}$ $\equiv\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }S_{g}c_{g}\leq 0\text{ and }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},$
[…]
Note that this formulation nests the case where only Assumption 1 holds that was presented in Proposition 1 by letting, for example, $S_{g}=0_{s_{g}\times K}$ the $s_{g}\times K$ zero matrix for all $g$. Thus in what follows I will analyze the program of Proposition 3 even when only Assumption 1 holds.
There is no longer a closed-form solution to the problem of
$\min_{\{c_{kg}\}\in[0,1]^{K}}\underset{\{c_{kg}\}\in[0,1]^{K}}{(\max)}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }S_{g}c_{g}\leq 0\text{ and }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ (8)
given a particular $\{p_{kg}\}\in P_{g}$, as there was in Corollary 1. However, given $\{p_{kg}\}\in P_{g}$ the solution to (8) is a linear program. Thus we can again solve for $L_{g}$ and $U_{g}$ by searching for the minimum and maximum respectively of the linear program (8) over all $\{p_{kg}\}\in P_{g}$. The suggested method of using a nonconvex solver with a coarse grid of starting points to solve the problem is still valid.
Remark 2.
This methodology is flexible and can easily be adjusted to accommodate further assumptions on the underlying individual-level model through addition restrictions to the optimization problem. For example, if some of the covariances between covariates can be estimated or bounded using a separate data set, this could be incorporated as polyhedral restrictions on the joint support of the covariates in the $P_{g}$ optimization problem. In addition, restrictions on the underlying distribution of $X_{i}$ can be incorporated through specification of the support.
Remark 3.
If we impose the assumption that the $Y_{i},X_{1i},\dots,X_{Li}$ were mutually independent conditional on $G_{i}$ we would obtain point identification of $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ and thus point identification of $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$. If we only impose that the $X_{1i},\dots,X_{Li}$ were mutually independent conditional on $G_{i}$ we would obtain point identification of the joint distribution of $X_{i}|G_{i}$. Sharp bounds on $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ follow from Cross and Manski (2002): we can consider the $L_{g}$ and $U_{g}$ optimization problems with $p_{kg}$ the joint distribution of $X_{i}$.
3 Estimation and Inference
3.1 Estimation
In practice, we observe sample analogs of the population values $\mathbb{E}[Y_{i}|G_{i}=g],\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g],\mathbb{P}[G_{i}=g]$ in the aggregate data set. For all $\ell=1,\dots,L,j=1,\dots,K,g=1,\dots,G$, denote
$\bar{Y}_{g}$ $=\frac{1}{n}\sum_{i=1}^{n}Y_{i}\mathbbm{1}\{G_{i}=g\}$ (9)
$\widehat{Pr}[X_{\ell i}=x_{j,\ell}|G_{i}=g]$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{X_{\ell i}=x_{j,\ell}\}\mathbbm{1}\{G_{i}=g\}$ (10)
$\widehat{Pr}[G_{i}=g]$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{G_{i}=g\}.$ (11)
$\bar{Y}_{g},\widehat{Pr}[X_{\ell i}=x_{j,\ell}],$ and $\widehat{Pr}[G_{i}=g]$ all converge to the respective population values by the law of large numbers. Thus we can construct a plug-in estimator, denote $\hat{D}$, for the sharp identified set by replacing all population values in the optimization problems of Proposition 3 with their sample estimates:
$\hat{D}\equiv\left[\hat{L},\hat{U}\right]\equiv\left[\sum_{g=1}^{G}\widehat{Pr}[G_{i}=g]\hat{L}_{g},\sum_{g=1}^{G}\widehat{Pr}[G_{i}=g]\hat{U}_{g}\right],$ (12)
$\hat{L}_{g}$ $\equiv\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }S_{g}c_{g}\leq 0~{}\forall g\text{ and }\exists\{p_{kg}\}\in\hat{P}_{g}\text{ with }\bar{Y}_{g}=\sum_{k=1}^{K}c_{kg}p_{kg},$ (13)
$\hat{U}_{g}$ $\equiv\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }S_{g}c_{g}\leq 0~{}\forall g\text{ and }\exists\{p_{kg}\}\in\hat{P}_{g}\text{ with }\bar{Y}_{g}=\sum_{k=1}^{K}c_{kg}p_{kg},$ (14)
$\hat{P}_{g}\equiv\text{arg}\min_{\{p_{kg}\}}\sum_{r=1}^{LK}v_{r}^{+}+v_{r}^{-}\;\;\text{s.t.}\sum_{k=1}^{K}p_{kg}=1;p_{kg},v_{r}^{+},v_{r}^{-}\geq 0~{}\forall k,r;\text{ and }\\[-10.0pt] \widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]-\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}=v_{K(\ell-1)+k}^{+}-v_{K(\ell-1)+k}^{-}~{}\forall\ell,k.$ (15)
The following proposition shows that the lower and upper bounds of the plug-in estimated set $\hat{D}$ are consistent.
Proposition 4.
[…]
Remark 4.
A similar plug-in estimator for the bounds provided in Proposition 2 from the Fréchet inequalities can be constructed. These estimated lower and upper bounds will be consistent for the lower and upper bounds of the identified set of Proposition 2 by the continuous mapping theorem.
3.2 Inference
Existing inference methods for partially identified estimated sets usually require knowledge of the joint distribution of the individual-level data, to estimate a covariance matrix used in constructing critical values or test statistics for valid coverage. However, in my setting I only observe marginal distributions of each variable. I point out why existing methods cannot be applied for a few, by no means representative, examples below:
Example.
Horowitz and Manski (2000) derive analytic bounds on conditional mean outcomes and point out that the delta method delivers asymptotic normality of the lower and upper bound estimators. The paper bootstraps the asymptotic covariance matrix to obtain a confidence interval that contains the identified set with correct asymptotic coverage. Putting aside that I do not have analytic bounds in my setting, being able to derive the asymptotic covariance of bounds $\hat{L}_{g}$ and $\hat{U}_{g}$ requires that I know the covariances between, for example, $Y_{i}$ and any $X_{\ell i}$ or any $X_{{\ell_{1}}i}$ and $X_{{\ell_{2}}i}$ given $G_{i}$. However I only observe sample marginal distributions of each variable and thus cannot hope to estimate the covariance matrix. Bootstrapping will also not be possible because I do not observe the individual-level data and so cannot generate a bootstrap sample.
Example.
Imbens and Manski (2004) provide a method for constructing confidence intervals on the parameter value of interest instead of on the entire identified set. This method chooses critical values for correct coverage by again relying on joint asymptotic normality of the lower and upper bound estimators, $\hat{L}_{g}$ and $\hat{U}_{g}$ in my setting. As discussed in the above example, I cannot hope to estimate the variances of the lower and upper bounds with the marginal information observed in aggregate data alone.
Example.
Hsieh et al. (2022) construct confidence intervals for identified sets of solutions to convex optimization programs, specifically linear and quadratic optimization programs with estimated coefficients, exploiting the necessary and sufficient optimality conditions. Putting aside that the optimization program is nonconvex in my setting, implementation of this inference method requires the asymptotic covariance matrix of the estimated covariates of the optimization problem, which in my setting rely on $\bar{Y}_{g},\widehat{Pr}[X_{\ell i}=x_{j,\ell}|G_{i}=g],$ and $\widehat{Pr}[G_{i}=g]$. Thus I again need to know the covariances between, for example, $Y_{i}$ and any $X_{\ell i}$ or any $X_{{\ell_{1}}i}$ and $X_{{\ell_{2}}i}$ given $G_{i}$.
Therefore in this setting, I am forced to rely on inference methods that require only marginal information of each variable. I choose to use the Bonferroni correction to make marginal confidence intervals on each sample observation jointly valid across the whole sample. On the other hand, note that each aggregate observation is the sample average of a binary random variable, as can be seen from equations (9), (10), and (11). Thus I can use finite-sample valid binomial proportion confidence intervals for each sample observation, instead of relying on normal approximations. I choose to use Clopper-Pearson intervals to construct marginal confidence intervals on each observation.
Suppose we also observe $n$ in the data set, where $n$ is the number of individuals across all groups over which we aggregate to obtain the aggregate data. Let $M$ be the total number of observations in the aggregate data set, that is, the total number of $\bar{Y}_{g},\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g],$ and $\widehat{Pr}[G_{i}=g]$ observations in the data set across all groups $g$, support points $k$ and covariates $\ell$.
The inference procedure is as follows:
  1. 1.
    For every sample observation $\hat{p}$, construct two-sided level $1-\frac{\alpha}{M}$ Clopper-Pearson CIs, denoted $\left[p_{L},p_{U}\right]$. For $X=N\hat{p}$, the Clopper-Pearson CI is determined by quantiles of the beta distribution:
    $\left[p_{L},p_{U}\right]=\left[B\bigg{(}\frac{\alpha}{2M},X,N-X+1\bigg{)},B\bigg{(}1-\frac{\alpha}{2M},X+1,N-X\bigg{)}\right]$
    The resulting confidence intervals are $\left[\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]_{L},\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]_{U}\right]$ for each $\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$, $\left[\bar{Y}_{g,L},\bar{Y}_{g,U}\right]$ for each $\bar{Y}_{g}$, and $\left[\widehat{Pr}[G_{i}=g]_{L},\widehat{Pr}[G_{i}=g]_{U}\right]$ for each $\widehat{Pr}[G_{i}=g]$.
  2. 2.
    Solve the optimization programs of (13), (14), and (15) for all values of sample observations within the marginal confidence intervals constructed in step 1:
    $a)\quad\hat{P}_{g,CI}\equiv\text{arg}\min_{\{p_{kg}\}}\sum_{r=1}^{2LK}v_{r}^{+}+v_{r}^{-}\;\;\text{s.t.}\sum_{k=1}^{K}p_{kg}=$  $\;1;p_{kg},v_{r}^{+},v_{r}^{-}\geq 0~{}\forall k,r;$
    $\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]_{L}-\sum_{k=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}\leq$ $\;v_{K(\ell-1)+k}^{+}-v_{K(\ell-1)+k}^{-}~{}\forall\ell,k,\text{ and }$
    $\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]_{U}-\sum_{k=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}$ $\geq v_{K(L+\ell-1)+k}^{+}-v_{K(L+\ell-1)+k}^{-}~{}\forall\ell,k.$
    $b)\quad\hat{L}_{g,CI}\equiv\min_{\{c_{gk}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{gk}\text{ s.t. }S_{g}c_{g}\leq 0~{}\forall g\text{ and }\exists\{p_{gk}\}\in\hat{P}_{g,CI}\\[-11.0pt] \text{with }\bar{Y}_{g,L}\leq\sum_{k=1}^{K}c_{gk}p_{gk}\text{ and }\bar{Y}_{g,U}\geq\sum_{k=1}^{K}c_{gk}p_{gk},$
    $c)\quad\hat{U}_{g,CI}\equiv\max_{\{c_{gk}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{gk}\text{ s.t. }S_{g}c_{g}\leq 0~{}\forall g\text{ and }\exists\{p_{gk}\}\in\hat{P}_{g,CI}\\[-11.0pt] \text{with }\bar{Y}_{g,L}\leq\sum_{k=1}^{K}c_{gk}p_{gk}\text{ and }\bar{Y}_{g,U}\geq\sum_{k=1}^{K}c_{gk}p_{gk}.\vspace{0pt}$
  3. 3.
    The confidence interval is given by
    $\hat{D}_{CI}\equiv\left[\sum_{g=1}^{G}\widehat{Pr}[G_{i}=g]_{L}\hat{L}_{g,CI},\sum_{g=1}^{G}\widehat{Pr}[G_{i}=g]_{U}\hat{U}_{g,CI}\right].$
Proposition 5.
Suppose Assumptions 1 and 2 hold. Suppose also that the $\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ are valid marginal distributions for $X_{i}$ with respect to the assumed support. Then $\mathbb{P}[D\subseteq\hat{D}_{CI}]\geq 1-\alpha$.
Since $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]\in D$, this means that $\mathbb{P}[\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]\in\hat{D}_{CI}]\geq 1-\alpha$, as discussed in Imbens and Manski (2004). Proposition 5 says that the confidence interval $\hat{D}_{CI}$ has correct coverage for the identified set and thus for the identified parameter.
Remark 5.
To perform inference on the bounds from Proposition 2 from the Fréchet inequalities, a similar approach can be used. If jointly valid marginal confidence intervals are constructed on each sample observation, calculating the estimated bounds but plugging lower confidence interval bounds into any lower bounds and upper confidence interval bounds into any upper bounds will produce a confidence interval with correct coverage.
4 Simulations and Empirical Application
One setting in which publicly available data is in aggregate form is standardized exam data. In this section I will apply the methodology developed in the previous sections to construct bounds on conditional exam pass rates. In this application I focus on exam pass rates for English and math exams and student demographic information for the state of Rhode Island in spring of 2018 and spring of 2019.888This data is a subset of the data used by Jack et al. (2023) in their paper studying the effect of online learning during the COVID-19 pandemic on exam results; I thank the authors for their assistance with obtaining the data.
In my application I produce sharp bounds on the average pass rate conditional on three covariates: race (white versus non-white), eligibility for free and reduced price lunch (FRPL), and English-language learner (ELL) status. In Section 4.1 I first explore what causes the width of the bounds to vary in simulations calibrated to the Rhode Island aggregate data. In Section 4.2 I then present the empirical application, where I estimate bounds with and without various monotonicity shape restrictions.
4.1 Simulations
I present three different simulation exercises. In all exercises I have three binary covariates $white_{i},FRPL_{i}$, and $ELL_{i}$, and a binary outcome $pass_{i}$. There are 55 aggregate groups in each example, where in the Rhode Island data a group is a school district-year pair (each school district in a given year), and in all simulation exercises I assume there are 2000 individuals in each group.
In the first exercises, I choose a joint distribution such that the marginal distribution over groups of each covariate approximately matches the marginal distribution over groups of each aggregate-level covariate in the Rhode Island data, as plotted in Figure 1.
Figure 1: Sample distributions of aggregate variables in simulation exercise 1
Table 1 presents results from the first simulation exercise. Estimated bounds in Column 2 are incredibly wide, with most of them being uninformative (equal to $[0,1]$). The 95% confidence intervals presented in Column 3 are also wide, but not too much wider than the estimated bounds themselves. The parameters for which I obtain informative bounds seem to be those where the conditioning population is well-represented in the data: the simulated data has lots of groups with a high fraction of white students and low fractions of FRPL and ELL students, and the parameter for which I obtain the tightest bounds is the average pass rate among white, non-FRPL, non-ELL students. Sharp bounds on the difference between parameters are very similar to the Minkowski set difference between bounds on each of the parameters. In particular, as noted in Remark 1 the bounds contain 0. In fact, Remark 1 states that all bounds on the difference between parameters will contain 0 as I do not impose additional assumptions.
Table 1: Estimated bounds on conditional pass rate
True Value Estimated Bounds 95% CI Bounds on Difference
Parameter (1) (2) (3) (4)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ 0.618 $[0.162,0.863]$ $[0.156,0.877]$ $[-0.838,0.863]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$ 0.5 $[0,1]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ 0.274 $[0,1]$ $[0,1]$ $[-0.961,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ 0.184 $[0,0.961]$ $[0,0.966]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ 0.460 $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ 0.345 $[0,1]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ 0.159 $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ 0.097 $[0,1]$ $[0,1]$
Notes: 95% CIs are confidence intervals on the estimated bounds in column 2. CIs are constructed using the method from Section 3.2, taking $\mathbb{P}[G_{i}=g]$ as observed instead of $\widehat{Pr}[G_{i}=g]$. Bounds on the difference are sharp bounds on the top minus the bottom parameter for which each set of bounds are reported.
To see if the estimated bounds could be narrower, in the second simulation exercise I choose a joint distribution that produces the same true conditional average pass rates, but such that the marginal distribution over groups of each covariate is even closer to either 0 or 1, as can be seen in Figure 2. Results are presented in Table 2. In Column 2, bounds on the parameter for which the conditioning population is well-represented in the data are narrower and more informative than in the first exercise, but bounds on all other parameters are uninformative, while in the first exercise some of the bounds on other parameters were informative. However, bounds on the difference in the first row of Column 4 are still wide, even though they are narrower than in the first exercise. This is likely because while there is more information about the first parameter, there is less information on the second parameter so that bounds on the difference are still wide.
Figure 2: Sample distributions of aggregate variables in simulation exercise 2
Table 2: Estimated bounds on conditional pass rate
True Value Estimated Bounds 95% CI Bounds on Difference
Parameter (1) (2) (3) (4)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ 0.618 $[0.568,0.749]$ $[0.561,0.760]$ $[-0.432,0.749]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$ 0.5 $[0,1]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ 0.274 $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ 0.184 $[0,1]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ 0.460 $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ 0.345 $[0,1]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ 0.159 $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ 0.097 $[0,1]$ $[0,1]$
Notes: 95% CIs are confidence intervals on the estimated bounds in column 2. CIs are constructed using the method from Section 3.2, taking $\mathbb{P}[G_{i}=g]$ as observed instead of $\widehat{Pr}[G_{i}=g]$. Bounds on the difference are sharp bounds on the top minus the bottom parameter for which each set of bounds are reported.
This suggests that if the conditioning populations for the first two parameters were well-represented, bounds on the difference might be narrower. In the third simulation exercise I choose a joint distribution that produces the same true conditional average pass rates and marginal distributions over groups for $\mathbb{E}[FRPL_{i}]$ and $\mathbb{E}[ELL_{i}]$, but I let some groups have $\mathbb{E}[white_{i}]$ close to 0 and others have $\mathbb{E}[white_{i}]$ close to 1, as can be seen in Figure 3. Results are presented in Table 3. As expected, bounds on each parameter in Column 3 are relatively narrow on the parameters for which the conditioning population is well-represented in the data. Notably relative to the second exercise, bounds on the first parameter are wider and bounds on the second parameter are narrower. This suggests that there is a trade-off between obtaining narrow bounds on a single parameter and obtaining narrow bounds on multiple parameters. Loosely, the groups with $\mathbb{E}[white_{i}]$ close to 0 help to make bounds on the second parameter narrow but make bounds on the first parameter wider.
Figure 3: Sample distributions of aggregate variables in simulation exercise 3
The bounds on the difference in the first row of Column 4 are narrower than in the first or the second exercise, but are still wide and, important to signing the parameter, contain 0. This means that, from this simulated data set, it would not be possible to say whether the white/non-white average pass rate gap among any group of FRPL/non-FRPL and ELL/non-ELL students is positive or negative. This suggests that without further restrictions, obtaining usefully informative bounds on average marginal effects is challenging.
Table 3: Estimated bounds on conditional pass rate
True Value Estimated Bounds 95% CI Bounds on Difference
Parameter (1) (2) (3) (4)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ 0.618 $[0.337,0.857]$ $[0.333,0.862]$ $[-0.212,0.857]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$ 0.5 $[0,0.549]$ $[0,0.550]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ 0.274 $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ 0.184 $[0,1]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ 0.460 $[0,1]$ $[0,1]$ $[-0.912,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ 0.345 $[0,0.912]$ $[0,0.935]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ 0.159 $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ 0.097 $[0,1]$ $[0,1]$
Notes: 95% CIs are confidence intervals on the estimated bounds in column 2. CIs are constructed using the method from Section 3.2, taking $\mathbb{P}[G_{i}=g]$ as observed instead of $\widehat{Pr}[G_{i}=g]$. Bounds on the difference are sharp bounds on the top minus the bottom parameter for which each set of bounds are reported.
4.2 Empirical application
The results of the simulation exercise suggest that bounds on white/non-white average pass rate gaps will not be informative in the Rhode Island data. I thus consider imposing several monotonicity shape restrictions to see whether such shape restrictions can help to make the bounds narrower.
Motivated by test score gaps that have been documented between rich and poor students (Tavernise, 2012; Porter, 2015), I impose that for each value of $(white_{i},ELL_{i})$ the average pass rate is lower for FRPL students than for non-FRPL students:
$\mathbb{E}[pass_{i}|FRPL_{i}=1,white_{i},ELL_{i}]-\mathbb{E}[pass_{i}|FRPL_{i}=0,white_{i},ELL_{i}]\leq 0.$ (16)
I first present estimated bounds on math exam white/non-white average pass rate gaps in Table 4. As expected, sharp bounds on the white/non-white average pass rate gaps reported in Column 1 are wide and either uninformative or close to uninformative. Imposing the additional monotonicity restriction (16) helps narrow the bounds for some parameters, reported in Column 2, but bounds are still wide and contain 0.
Table 4: Rhode Island white/non-white math exam pass rate differences
Bounds without Monotonicity Monotonicity Bounds
Parameter (1) (2)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.940,0.850]$ $[-0.872,0.850]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-0.821,0.994]$ $[-0.691,0.596]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-1,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-1,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$
Notes: Bounds on difference impose no shape restrictions; monotonicity bounds impose monotonicity restriction (16), as discussed in Section 4.2.
For English exams, I impose an additional monotonicity restriction that for each value of $(white_{i},FRPL_{i})$ the average pass rate is lower for English-language learner students than for non-English-language learner students:
$\mathbb{E}[pass_{i}|ELL_{i}=1,white_{i},FRPL_{i}]-\mathbb{E}[pass_{i}|ELL_{i}=0,white_{i},FRPL_{i}]\leq 0.$ (17)
I present estimated bounds on English exam white/non-white average pass rate gaps in Table 5. Sharp bounds on the white/non-white average pass rate gaps reported in Column 1 are again wide and either uninformative or close to uninformative. Imposing both additional monotonicity restrictions of (16) and (17) helps to narrow the bounds all parameters, reported in Column 2, but all bounds are still wide and contain 0. Thus even with additional monotonicity restrictions, it is difficult to obtain bounds that are narrow enough to be useful from realistic aggregate data alone.
Table 5: Rhode Island white/non-white English exam pass rate differences
Bounds without Monotonicity Monotonicity Bounds
Parameter (1) (2)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.904,0.923]$ $[-0.823,0.923]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-0.852,1]$ $[-0.759,0.706]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-1,0.914]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-0.759,0.706]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$
Notes: Bounds on difference impose no shape restrictions; monotonicity bounds impose monotonicity restrictions (16) and (17), as discussed in Section 4.2.
5 Conclusion
In this paper I present sharp bounds on individual-level parameters of interest when the only available data is aggregate data, and I develop a valid inference method relying only on the available marginal information of each covariate. In simulations and an empirical application I show that the sharp bounds are too wide to be useful with realistic aggregate data, even after imposing additional shape restrictions. This suggests that researchers hoping to use aggregate data to make individual-level claims should either obtain additional individual-level data, or impose stronger assumptions to be able to use other existing methodology. However because the sharp bounds are so wide, stronger assumptions that deliver narrower bounds will have strong implications about the underlying aggregation process, meaning the researcher should always carefully justify any imposed assumptions.
Appendix A Appendix: Proofs
A.1 Proof of Lemma 1
Proof.
As argued in the main text, the only information we have in addition to Assumption 1 are equations (1), (2), (3), and that $\delta_{k},\gamma_{kg},\pi_{kg}\in[0,1]$ for all $k,g$. The set given by (4) imposes all of these restrictions and nothing more, and finds the set of $\sum_{k=1}^{K}\lambda_{k}\delta_{k}$ such that the restrictions are satisfied. Thus the set is sharp. ∎
A.2 Proof of Proposition 1
Proof.
For any $g$, given any $\{p_{kg}\}_{k=1}^{K}$ such that $\sum_{k=1}^{K}p_{kg}=1$ and $p_{kg}\geq 0~{}\forall k$, note that there exists $\{c_{kg}\}_{k=1}^{K}\in[0,1]^{K}$ such that $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$; namely $c_{kg}=\mathbb{E}[Y_{i}|G_{i}=g]~{}\forall k$.
Thus imposing the restriction $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ does not further restrict the set of possible $\{p_{kg}\}_{k=1}^{K}$ if the following restrictions already hold: $\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}$, $\sum_{k=1}^{K}p_{kg}=1$ and $p_{kg}\geq 0~{}\forall k$. Finally note that all restrictions involving $\{p_{kg}\}_{k=1}^{K}$ only involve $p_{kg}$ with the same index $g$. This means it is equivalent to write (4) as
$D=\bigg{\{}\sum_{k=1}^{K}\lambda_{k}d_{k}\bigg{|}~{}0\leq d_{k}\leq 1~{}\forall k,\text{ and }\exists\,(p_{1g},\dots,p_{Kg})\in P_{g},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }d_{k}=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}~{}\forall k,~{}\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g\bigg{\}},$ (A.1)
where
$P_{g}=\bigg{\{}(p_{1g},\dots,p_{Kg})\in[0,1]^{K}\bigg{|}~{}\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}~{}\forall\ell,k,\text{ and }\sum_{k=1}^{K}p_{kg}=1\bigg{\}}.$ (A.2)
If $\mathbb{P}[G_{i}=g],c_{kg}\in[0,1]$ for all $k$ and $g$ then because $d_{k}=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}$ we know $d_{k}\in[0,1]$ for all $k$. Then we can get rid of $d_{k}$ in (A.1) by plugging in one of the constraints like so:
$D=\bigg{\{}\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}\bigg{|}~{}\exists\,(p_{1g},\dots,p_{Kg})\in P_{g},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g\bigg{\}}.$ (A.3)
I next show that $D$ is an interval. For any given $\{p_{kg}\}\in P_{g}$, we argued above that $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\mathbb{E}[Y_{i}|G_{i}=g]\in D$. For any $g$ consider arbitary $\{c_{kg}\},\{c_{kg}^{\prime}\}\in[0,1]^{K}$ with corresponding $\{p_{kg}\},\{p_{kg}^{\prime}\}\in P_{g}$ such that $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ and $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}^{\prime}p_{kg}^{\prime}$. We know $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}\in D$ and $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}^{\prime}\in D$. For arbitrary $t\in[0,1]$, let
$\tilde{c}_{kg}=tc_{kg}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$
for all $k,g$ and let
$\tilde{c}_{kg}^{\prime}=tc_{kg}^{\prime}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$
for all $k,g$. Then $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k}p_{kg}\tilde{c}_{kg}=\sum_{k}p_{kg}^{\prime}\tilde{c}_{kg}^{\prime}$ for all $g$ and $\tilde{c}_{kg},\tilde{c}_{kg}^{\prime}\in[0,1]$ for all $k,g$. Thus $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\tilde{c}_{kg}\in D$ and $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\tilde{c}_{kg}^{\prime}\in D$. Note
$\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\tilde{c}_{kg}$
$=t\left(\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}\right)+(1-t)\left(\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\mathbb{E}[Y_{i}|G_{i}=g]\right),$
$\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\tilde{c}_{kg}^{\prime}$
$=t\left(\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}^{\prime}\right)+(1-t)\left(\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\mathbb{E}[Y_{i}|G_{i}=g]\right).$
Since $t$ was arbitrary between 0 and 1, $D$ contains any value between $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\mathbb{E}[Y_{i}|G_{i}=g]$. Similarly, $D$ contains any value between $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}^{\prime}$ and $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\mathbb{E}[Y_{i}|G_{i}=g]$. In particular, $D$ contains any value between $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}^{\prime}$. Thus $D$ is an interval.
This means that we can equivalently express (A.3) and (A.2) as optimization problems as follows: $D=\left[L,U\right]$, where
$L$ $\equiv\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},$
$U$ $\equiv\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},\text{ and}$
$P_{g}=\underset{\{p_{kg}\}\in[0,1]^{K}}{\text{arg}\min}\sum_{r=1}^{LK}v_{r}^{+}+v_{r}^{-}\;\;\text{s.t. }\sum_{k=1}^{K}p_{kg}=1,~{}p_{kg},v_{r}^{+},v_{r}^{-}\geq 0~{}\forall k,r,\text{ and}\\[-10.0pt] \mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]-\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}=v_{K(\ell-1)+k}^{+}-v_{K(\ell-1)+k}^{-}~{}\forall\ell,k.$
The slack variables in the $P_{g}$ problem will all be equal to zero at optimum.
Finally, since the constraints in $L$ and $U$ are for each $g$, we can equivalently solve a program for each group $g$ and take a weighted sum of the solutions to get lower and upper bounds for $D$, like so: $D=\left[\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]L_{g},\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]U_{g}\right]$, where
$L_{g}$ $\equiv\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},$
$U_{g}$ $\equiv\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}.$
A.3 Proof of Corollary 1
Proof.
For given weights $\lambda_{1},\dots,\lambda_{K}$ and any fixed $\{p_{kg}\}\in P_{g}$, relabel the indices $k=1,\dots,K$ so that $\frac{\lambda_{1}}{p_{1g}}\geq\dots\geq\frac{\lambda_{K}}{p_{Kg}}$, where if $p_{kg}=0$ we define $\frac{\lambda_{k}}{p_{gk}}\equiv+\infty$.
Note that the proposed solution in Corollary 1 is feasible. We will take advantage of strong duality (because Slater’s condition holds) and use the joint feasiblity and satisfaction of complementary slackness for proposed solutions to the primal and dual problems to show optimality.
The dual of the minimization linear program is
$\max_{u\in\mathbb{R},v\in\mathbb{R}^{2K}}-\mathbb{E}[Y_{i}|G_{i}=g]u-\sum_{i=1}^{K}v_{i}\text{ s.t. }-p_{kg}u-v_{k}\leq\lambda_{k}~{}\forall k=1,\dots,K,v_{i}\geq 0~{}\forall i=1,\dots,K.$
Let $k$ be such that $\sum_{j=1}^{k-1}p_{jg}\leq 1-\mathbb{E}[Y_{i}|G_{i}=g]<\sum_{j=1}^{k}p_{jg}$ holds. Consider a solution to the dual where $v_{i}=0$ for all $i\leq k$, and for $i\geq k$ we have that $v_{i}$ satisfies $-p_{ig}u-v_{i}=\lambda_{i}$, meaning $u=-\frac{\lambda_{k}}{p_{kg}}$. Thus for $i>k$ we have $v_{i}>0$ because $-p_{ig}u=\frac{\lambda_{k}p_{ig}}{p_{kg}}\geq\lambda_{i}$. Clearly this is a feasible solution.
We see that complementary slackness holds with the condition that $c_{ig}\leq 1$ for all $i$ and $v_{i}\geq 0$ for all $i$ because $c_{ig}\neq 1$ for all $i\leq k$ while $v_{i}=0$ for all $i\leq k$. Complementary slackness holds with the condition that $c_{ig}\geq 0$ for all $i$ and $-p_{ig}u-v_{i}\leq\lambda_{i}$ for all $i$ because $c_{ig}\neq 0$ for all $i\geq k$ while $-p_{ig}u-v_{i}=\lambda_{i}$ for all $i\geq k$.
Thus the proposed solution for the minimization problem is optimal.
The dual of the maximization linear program is
$\min_{u\in\mathbb{R},v\in\mathbb{R}^{2K}}\mathbb{E}[Y_{i}|G_{i}=g]u+\sum_{i=1}^{K}v_{i}\text{ s.t. }p_{k}u+v_{k}\geq\lambda_{k}~{}\forall k=1,\dots,K,v_{i}\geq 0~{}\forall i=1,\dots,K.$
Let $k$ be such that $\sum_{j=1}^{k-1}p_{jg}\leq\mathbb{E}[Y_{i}|G_{i}=g]<\sum_{j=1}^{k}p_{jg}$ holds. Consider a solution to the dual where $v_{i}=0$ for all $i\geq k$, and for $i\leq k$ we have that $v_{i}$ satisfies $p_{ig}u+v_{i}=\lambda_{i}$, meaning $u=\frac{\lambda_{k}}{p_{kg}}$. Thus for $i<k$ we have $v_{i}>0$ because $p_{ig}u=\frac{\lambda_{k}p_{ig}}{p_{kg}}\leq\lambda_{i}$. Clearly this is a feasible solution.
We see that complementary slackness holds with the condition that $c_{ig}\leq 1$ for all $i$ and $v_{i}\geq 0$ for all $i$ because $c_{ig}\neq 1$ for all $i\geq k$ while $v_{i}=0$ for all $i\geq k$. Complementary slackness holds with the condition that $c_{ig}\geq 0$ for all $i$ and $p_{ig}u+v_{i}\geq\lambda_{i}$ for all $i$ because $c_{ig}\neq 0$ for all $i\leq k$ while $p_{ig}u+v_{i}=\lambda_{i}$ for all $i\leq k$.
Thus the proposed solution for the minimization problem is optimal. ∎
A.4 Proof of Proposition 2
Proof.
First I show that $D\subseteq D^{F}$.
Since $D$ is truly an interval, as proved when proving Proposition 1, there exists some $\{c_{kg}\}_{k,g},\{p_{kg}\}_{k,g}$ that satisfy the constraints of $L_{g},U_{g},$ and $P_{g}$ where $\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\sum_{k=1}^{K}\lambda_{k}c_{kg}\in D$.
By an argument analogous to that given in the proof of Proposition 1, the set of each $c_{kg}$ that satisfy the constraints is the sharp identified set for $\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]$. Thus it is sufficient to show that for each $k,g$,
$\frac{L_{g}(1,x_{k})}{L_{g}(1,x_{k})+U_{g}(0,x_{k})}\leq\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]\leq\frac{U_{g}(1,x_{k})}{U_{g}(1,x_{k})+L_{g}(0,x_{k})}.$
This means $\left[\frac{L_{g}(1,x_{k})}{L_{g}(1,x_{k})+U_{g}(0,x_{k})},\frac{U_{g}(1,x_{k})}{U_{g}(1,x_{k})+L_{g}(0,x_{k})}\right]$ are also bounds, so it follows that
$\frac{L_{g}(1,x_{k})}{L_{g}(1,x_{k})+U_{g}(0,x_{k})}\leq c_{kg}\leq\frac{U_{g}(1,x_{k})}{U_{g}(1,x_{k})+L_{g}(0,x_{k})}.$
and thus $\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\sum_{k=1}^{K}\lambda_{k}c_{kg}\in D^{F}$.
Note
$\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]$ $=\mathbb{P}[Y_{i}=1|X_{i}=x_{k},G_{i}=g]$
$=\frac{\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]}{\mathbb{P}[X_{i}=x_{k}|G_{i}=g]}$
$=\frac{\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]}{\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]+\mathbb{P}[Y_{i}=0,X_{i}=x_{k}|G_{i}=g]}.$
One can check that the function $\frac{x}{x+y}$ is increasing in $x$ and decreasing in $y$; thus $\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]$ attains its minimum when $\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]$ is as small as possible and $\mathbb{P}[Y_{i}=0,X_{i}=x_{k}|G_{i}=g]$ is as large as possible. Similarly $\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]$ attains its maximum when $\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]$ is as large as possible and $\mathbb{P}[Y_{i}=0,X_{i}=x_{k}|G_{i}=g]$ is as small as possible.
Since $\mathbb{P}[Y_{i}=0,X_{i}=x_{k}|G_{i}=g]\in[L_{g}(0,x_{k}),U_{g}(0,x_{k})]$ and $\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]\in[L_{g}(1,x_{k}),U_{g}(1,x_{k})]$, it follows then that $\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]\in\left[\frac{L_{g}(1,x_{k})}{L_{g}(1,x_{k})+U_{g}(0,x_{k})},\frac{U_{g}(1,x_{k})}{U_{g}(1,x_{k})+L_{g}(0,x_{k})}\right]$.
Now note that if the Fréchet inequalities on $\mathbb{P}[Y_{i}=y,X_{i}=x_{k}|G_{i}=g]$ are sharp for all $y\in\{0,1\},k=1,\dots,K,g=1,\dots,G$ then $\mathbb{P}[Y_{i}=0,X_{i}=x_{k}|G_{i}=g]$ attains its minimum at $L_{g}(0,x_{k})$ and its maximum at $U_{g}(0,x_{k})$, and $\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]$ attains its minimum at $L_{g}(1,x_{k})$ and its maximum at $U_{g}(1,x_{k})$. Thus bounds $\left[\frac{L_{g}(1,x_{k})}{L_{g}(1,x_{k})+U_{g}(0,x_{k})},\frac{U_{g}(1,x_{k})}{U_{g}(1,x_{k})+L_{g}(0,x_{k})}\right]$ on $\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]$ are indeed sharp and so $D=D^{F}$. ∎
A.5 Proof of Proposition 3
Proof.
To prove sharpness of the identified set, it is sufficient to show that the following set is an interval: for $c_{g}\equiv(c_{1g},\dots,c_{Kg})^{\prime}$,
$\bigg{\{}\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}\bigg{|}~{}S_{g}c_{g}\leq 0~{}\forall g\text{ and }\exists\,(p_{1g},\dots,p_{Kg})\in P_{g},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g\bigg{\}}.$ (A.4)
We know that the set of $c_{g}\in[0,1]^{KG}$ such that $S_{g}c_{g}\leq 0$ for all $g$ is a convex, closed, and bounded set because it is the intersection of a closed, bounded, and convex set with a half-space. This means the set of values of $\sum_{k}\lambda_{k}c_{kg}$ such that $S_{g}c_{g}\leq 0$ for all $g$ is also a closed and bounded and convex set.
In the proof of Proposition 1 I argued that for any $c_{g},c_{g}^{\prime}\in[0,1]^{K}$ with corresponding $\{p_{kg}\},\{p_{kg}^{\prime}\}\in P_{g}$ that satisfy
$\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ $\qquad\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}^{\prime}p_{kg}^{\prime}$ (A.5)
$S_{g}c_{g}\leq 0$ $\qquad S_{g}c_{g}^{\prime}\leq 0,$ (A.6)
we know $tc_{kg}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$ and $tc_{kg}^{\prime}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$ also satisfy (A.5) in place of $c_{kg}$ and $c_{kg}^{\prime}$ respectively, for any $t\in[0,1]$. Thus for any $\sum_{k=1}^{K}\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}c_{kg}^{\prime}$, $\{\tilde{c}_{kg}\}$ satisfies $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}\tilde{c}_{kg}p_{kg}$. And that (A.6) holds implies $S_{g}(tc_{g}+(1-t)c_{g}^{\prime})\leq 0$ for any $t\in[0,1]$, meaning that for any $\sum_{k=1}^{K}\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}c_{kg}^{\prime}$, $\{\tilde{c}_{kg}\}$ satisfies $S_{g}\tilde{c}_{g}\leq 0$.
Thus any $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}^{\prime}$ is in the set given by (A.4), meaning (A.4) is an interval. Thus the sharp identified set is an interval. The rest of Proposition 3 can be proved following arguments similar to those used in the proof of Proposition 1. ∎
A.6 Proof of Proposition 4
Proof.
The result follows from the Theorem of the Maximum for the following correspondence, defining $c=(c_{1},\dots,c_{K})^{\prime}$:
$\Gamma\left(y,\{g_{k,\ell}\}_{k,\ell}\right)=\underset{\{p_{k},c_{k}\}_{k}}{\text{arg}\max}\left(\underset{\{p_{k},c_{k}\}_{k}}{\text{arg}\min}\right)\bigg{\{}\sum_{k=1}^{K}\lambda_{k}c_{k}\text{ s.t. }S_{g}c\leq 0,y=\sum_{k=1}^{K}c_{k}p_{k},\sum_{k=1}^{K}p_{k}=1,\\ \qquad g_{k,\ell}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j},c_{k}\in[0,1]~{}\forall k,p_{k}\geq 0~{}\forall k\bigg{\}}.$
Since the $\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ are valid marginal distributions for $X_{i}$ with respect to the assumed support, this correspondence is analogous to the set defined in Proposition 3; that is, the constraints in $P_{g}$ hold with $v_{r}^{+}=v_{r}^{-}=0$.
To apply the Theorem of the Maximum, I must show $\Gamma$ is a continuous and compact-valued correspondence. That $\Gamma$ is compact-valued follows because the constraints are intersections of non-parallel planes with a curve $(y=\sum_{k=1}^{K}c_{k}p_{k})$ on a compact set $(c_{k}\in[0,1],p_{k}\geq 0,\sum_{k=1}^{K}p_{k}=1)$.
Let $(y^{n},\{g_{k,\ell}^{n}\}_{k,\ell})$ be an arbitrary sequence such that $(y^{n},\{g_{k,\ell}^{n}\}_{\ell,j})\to(y,\{g_{k,\ell}\}_{k,\ell})$ as $n\to\infty$ and all $y^{n},y\in[0,1]$ and $\{g_{k,\ell}^{n}\}_{k,\ell}$ are valid marginal distributions given the assumed support. Let ($(\{p_{k}^{n}\},\{c_{k}^{n}\})$ be an arbitrary sequence such that ($(\{p_{k}^{n}\},\{c_{k}^{n}\})\to(\{p_{k}\},\{c_{k}\})$ as $n\to\infty$ and for each $n$, ($(\{p_{k}^{n}\},\{c_{k}^{n}\})\in\Gamma\left((y^{n},\{g_{k,\ell}^{n}\}_{k,\ell})\right)$. If we can show ($(\{p_{k}\},\{c_{k}\})\in\Gamma((y,\{g_{k,\ell}\}_{k,\ell}))$ then $\Gamma$ is upper hemicontinuous.
Since $\sum_{k=1}^{K}p_{k}^{n}=1$ for all $n$, $p_{k}^{n}\geq 0$ for all $k,n$, and $p_{k}^{n}\to p_{k}$ for all $k$, it follows that $\sum_{k=1}^{K}p_{k}=1$ and $p_{k}\geq 0$ for all $k$. Since $c_{k}^{n}\in[0,1]$ for all $k,n$ and $c_{k}^{n}\to c_{k}$ for all $k$, it follows that $c_{k}\in[0,1]$ for all $k$. Since $y^{n}=\sum_{k=1}^{K}p_{k}^{n}c_{k}^{n}$ for all $k,n$ and $y^{n}\to y,d_{k}^{n}\to d_{k},p_{k}^{n}\to p_{k}$ for all $k,n$, it follows that $y=\sum_{k}p_{k}d_{k}$. Since $g_{k,\ell}^{n}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}^{n}$ for all $k,\ell,n$ and we have $p_{k}^{n}\to p_{k},g_{k,\ell}^{n}\to g_{k,\ell}$ for all $k,\ell$, it follows that $g_{k,\ell}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}$ for all $k,\ell$. Finally since $S_{g}c^{n}\leq 0$ for all $n$ and we have $c^{n}\to c$ (this is shorthand for $\{c_{k}^{n}\}\to\{c_{k}\}$), it follows that $S_{g}c\leq 0$. Thus ($(\{p_{k}\},\{c_{k}\})\in\Gamma((y,\{g_{k,\ell}\}_{k,\ell}))$ and thus $\Gamma$ is upper hemicontinuous.
Let $(y^{n},\{g_{k,\ell}^{n}\}_{k,\ell})$ be an arbitrary sequence such that $(y^{n},\{g_{k,\ell}^{n}\}_{\ell,j})\to(y,\{g_{k,\ell}\}_{k,\ell})$ as $n\to\infty$ and all $y^{n},y\in[0,1]$ and $\{g_{k,\ell}^{n}\}_{k,\ell}$ are valid marginal distributions given the assumed support. Let ($(\{p_{k}\},\{c_{k}\})\in\Gamma(y,\{g_{k,\ell}\}_{k,\ell})$ be arbitrary. If we can show there exists a subsequence $(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$ and sequence ($(\{p_{k}^{t}\},\{c_{k}^{t}\})$ such that ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\to(\{p_{k}\},\{c_{k}\})$ as $t\to\infty$ and ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\in\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$ for all $t$, then $\Gamma$ is lower hemicontinuous.
We know that $g_{k,\ell}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}$ and $\sum_{k=1}^{K}p_{k}=1$. Because we assume each $\{g_{k,\ell}\}$ is a valid marginal distribution of covariates there are strictly less than $K$ unique noncollinear equations in this system of equations. We can express these equations as a linear system $Ap=g$, where $p=\{p_{k}\}$ and $g=\{g_{k,\ell}\}_{k,\ell}$, and all solutions are given by $p=A^{+}g+(I-A^{+}A)w$, where $A^{+}$ is the Moore-Penrose inverse and $w$ is any arbitrary vector of correct dimension.
Thus the solution $\{p_{k}\}$ to this system of linear equations is continuous in $\{g_{k,\ell}\}_{k,\ell}$. Therefore for each $t$ there exists $n_{t}$ large enough that we can find $\{p_{k}^{t}\}$ arbitrarily close to $\{p_{k}\}$ such that $g_{k,\ell}^{n_{t}}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}^{t}$ and $\sum_{k=1}^{K}p_{k}^{t}=1$. Such a solution exists because we assume $\hat{D}$ is always nonempty with a valid marginal distribution of covariates $\{g_{k,\ell}^{n_{t}}\}$.
Recall each $p_{k}\geq 0$. In fact, the proposition imposes as an assumption that we only need consider $\{p_{k}\}$ such that each $p_{k}>0$. Thus we can make $p_{k}^{t}\geq 0$ for each $k$. Thus we can construct a sequence $\{p_{k}^{t}\}$ with corresponding subsequence $\{g_{k,\ell}^{n_{t}}\}_{k,\ell}$ such that $g_{k,\ell}^{n_{t}}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}^{t}$ for all $t$, $\sum_{k=1}^{K}p_{k}^{t}=1$ for all $t$, $p_{k}^{t}\geq 0$ for all $k$, and $p_{k}^{t}\to p_{k}$ for all $k$ as $t\to\infty$.
Given this sequence $\{p_{k}^{t}\}$, we need to find a corresponding sequence $\{c_{k}^{t}\}$ and subsequence $\{y^{n_{t}}\}$ such that $y^{n_{t}}=\sum_{k=1}^{K}c_{k}^{t}p_{k}^{t}$ and $c_{k}^{t}\in[0,1]$ for all $k,t$. Note that we can write the equation $y=\sum_{k=1}^{K}c_{k}p_{k}$ as $p^{\prime}c=y$ and consider the Moore-Penrose representation of the set of all solutions $c$. As a matrix algebra result, it is true that the Moore-Penrose inverse of $p$ is continuous, that is $(p_{t})^{+}\to p^{+}$ as long as the rank of $p_{t}$ is the same as the rank of $p$ for all $t$. Since the rank of $p_{t}$ and $p$ is always 1, it follows that the solution $c$ to the equation is continuous in $p$ and $y$.
Therefore for each $t$ there exists $n_{t}$ large enough that we can find $\{\tilde{c}_{k}^{t}\}$ arbitrarily close to $\{c_{k}\}$ with $y^{n_{t}}=\sum_{k=1}^{K}\tilde{c}_{k}^{t}p_{k}^{t}$ (with some relabeling of the $\{p_{k}^{t}\}$ sequence indices as necessary). We know $c_{k}\in[0,1]$ for each $k$, but it may be the case that $\tilde{c}_{k}^{t}\notin[0,1]$ for some $k$. However for those $k$ we will have $\tilde{c}_{k}^{t}$ arbitrarily close to $[0,1]$. As argued in the proof of Proposition 1 it also holds that $y^{n_{t}}=\sum_{k=1}^{K}y^{n_{t}}p_{k}^{t}$ (and we know $y^{n,t}\in[0,1]$), meaning there exists another feasible $\{c_{k}^{t}\}$ between $\{\tilde{c}_{k}^{t}\}$ and $(y^{n_{t}},\dots,y^{n_{t}})$ that is arbitrarily close to $\{\tilde{c}_{k}^{t}\}$ but with $c_{k}^{t}\in[0,1]$ for all $k$.
Finally, we know that $S_{g}c\leq 0$. If $S_{g}c\leq 0$ imposes nontrivial restrictions on $c$, suppose that in fact $S_{g}c<0$. Since $S_{g}x$ is a continuous linear map in $x$ and we have $\{c_{k}^{t}\}\to c_{k}$, for large enough $t$ we have $S_{g}c^{t}\leq 0$ as well.
Thus we have sequence ($(\{p_{k}^{t}\},\{c_{k}^{t}\})$ with corresponding subsequence $(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$ such that ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\to(\{p_{k}\},\{c_{k}\})$ as $t\to\infty$, and ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\in\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$. Therefore $\Gamma$ is lower hemicontinuous.
[…]
Weak law of large numbers also gives us that $\widehat{Pr}[G_{i}=g]$ is consistent, so that by continuous mapping theorem again we have that the lower and upper bounds of $\hat{D}$ converge to the lower and upper bounds of $D$. ∎
A.7 Proof of Proposition 5
Proof.
Since Clopper-Pearson intervals are finite-sample valid for each sample observation, using the Bonferroni correction means that the population analog of every single sample observation is contained in its Clopper-Pearson interval with joint probability greater than $1-\alpha$.
Note that if all population analogs of the sample observations are jointly in their respective Clopper-Pearson intervals then any $\{c_{kg}\}$ and $\{p_{kg}\}$ consistent with the population $P_{g}$ and $L_{g},U_{g}$ constraints in the formulation of $D$ are also consistent with the $\hat{P}_{g,CI}$ and $\hat{L}_{g,CI},\hat{U}_{g,CI}$ constraints in the formulation of $\hat{D}_{CI}$. Thus the set of $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ consistent with the population $P_{g}$ and $L_{g},U_{g}$ constraints in the formulation of $D$ is a subset of the set of $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ consistent with the $\hat{P}_{g,CI}$ and $\hat{L}_{g,CI},\hat{U}_{g,CI}$ constraints in the formulation of $\hat{D}_{CI}$. And if the true $\mathbb{P}[G_{i}=g]$ are all contained in their Clopper-Pearson intervals, it follows that $D\subseteq\hat{D}_{CI}$.
Since the event that all population analogs of the sample observations are jointly in their respective Clopper-Pearson intervals happens with probability at least $1-\alpha$, it follows that $\mathbb{P}[D\subseteq\hat{D}_{CI}]\geq 1-\alpha$. ∎