### 05 Apr 2024 v3

Partial Identification of Individual-Level Parameters Using Aggregate Data in a Nonparametric Binary Outcome Modelthanks: I am especially grateful to Edward Vytlacil and Isaiah Andrews for their guidance and advice. I thank Alberto Abadie, Anna Mikusheva, Hyungsik Roger Moon, Elie Tamer, and participants in the MIT econometrics lunch seminar for their helpful comments. All errors are my own.
Sarah Moon Department of Economics, Massachusetts Institute of Technology, sarahmn@mit.edu
1 Introduction
Researchers frequently use publicly available data in policy analysis, which is usually provided at the aggregate level due to individual privacy concerns. For example, statewide standardized exam results are reported in the form of school-wide or school district-wide pass rates, as opposed to pass result and demographic information (e.g. gender, race, family income) for each individual student. Election data is available as vote shares and voter demographics over voting districts, as opposed to the vote and demographics information for each individual voter.
Many researchers use aggregate data to run analyses where the outcome of interest is at the individual level. One example is the literature on the relationship between religion and suicide, which often uses suicide rates and religion prevalence rates (Neeleman et al., 1997; Dervic et al., 2004; Becker and Woessmann, 2018). Jack et al. (2023) looks at the relationship between virtual learning during the COVID-19 pandemic and student performance in grade school using standardized exam pass rates and district demographic information. Alter et al. (1999) looks at the relationship between median neighborhood income, rate of use of cardiac procedures, and mortality rates after a heart attack to determine the relationship between patient income and treatment outcomes. Examples of aggregate data used in political science analyses of individual voting behavior are given in Freedman et al. (1998) and Kousser (2001). Another large literature studies the relationship between mortality and economic conditions and often utilizes aggregate data in the form of mortality rates and unemployment rates (Ruhm, 2000; Neumayer, 2004; Gerdtham and Ruhm, 2006; Svensson, 2007; Lindo, 2015).
It is well known that aggregate variables can be related in ways that are different from the same variables at the individual level, [^1] Firebaugh (1978) notes that “aggregate” and “individual” are relative terms, as individual variables correspond to the individual unit of analysis and need not correspond to an individual person. In the econometrics literature, the terms “macro” and “micro” are also often used. a problem known as an ecological fallacy. [^2] This problem has also been referred to as the problem of aggregation (Stoker, 1984). The ecological fallacy has been studied across many fields over the past several decades, beginning with the seminal work of Robinson (1950); Duncan and Davis (1953), and Theil (1954). These earlier papers demonstrated issues with interpreting parameters from regressions with aggregate data as individual-level effects, and provided simple bounds on individual-level effects with a binary outcome and single binary covariate. Since then, studies like Shiveley (1974); Kramer (1983); Greenland and Robins (1994); Gravelle et al. (2002); Hsiao et al. (2005), and Lindo (2015) have continued to demonstrate that results on individual-level effects are sensitive to the level of data aggregation and other imposed assumptions like functional form.
While the ecological fallacy has been acknowledged as an issue by the literature, to the best of my knowledge much of the literature is concerned with point identification of individual-level parameters (King et al., 1999; Rosen et al., 2001). [^3] One exception is Jiang et al. (2020), who develop partial identification techniques in a setting with two binary covariates, imposing linear individual-level relationships. When only aggregate information is available, knowledge about parameters at the individual level is limited, and point identification is often not achieved without further assumptions. Imposing more assumptions may allow for precise results, but such assumptions may be less plausible. For example, one popular method for analyzing aggregate data is the ecological inference method (King, 1997), used often in political science studies of elections (Burden and Kimball, 1998). This method relies on many assumptions, like imposing individual-level joint distributions and a lack of bias introduced by the aggregation, that often fail to hold in applied settings (Tam Cho, 1998; Cho and Gaines, 2004; Freedman et al., 1998; Kousser, 2001).
Even without these strong assumptions, partial identification is still possible. The resulting identified set is exactly the extent to which individual-level results are sensitive to assumptions. In this paper I consider the problem of identifying linear combinations of conditional mean outcomes, $\mathbb{E}[Y_{i}|X_{1i},\dots,X_{Li}]$, when individual-level outcome $Y_{i}$ is binary and the only data that is observed is the marginal distribution of each individual-level variable over many groups, which I call aggregate data. I develop a partial identification methodology that constructs sharp bounds by solving an optimization problem that considers all underlying joint distributions of individual-level covariates that are consistent with the observed marginal distributions. I show how further restrictions on the underlying data generating process, like shape restrictions or additional data at a finer level of aggregation, can be incorporated into the optimization problem to obtain sharp bounds. Since I do not observe individual-level joint distributions, I develop valid inference procedures on the identified set using marginal information only.
To demonstrate how informative these bounds can be, I apply this methodology to several different simulated aggregate data sets calibrated to the Rhode Island standardized exam score data set used in my empirical application. I find that bounds are relatively wide on marginal effect parameters of interest. Imposing monotonicity shape restrictions in the empirical application helps narrow bounds on test score gaps, and using additional data at a finer level of aggregation makes bounds on test score gaps even narrower. This suggests that it is easier to obtain useful individual-level results from aggregate data when the data is aggregated at a finer level and when the researcher can impose more restrictive assumptions about the individual-level data generating process.
Partial identification in various contexts has been widely studied in the econometrics literature, especially over the last thirty years. [^4] See Tamer (2010), Ho and Rosen (2015), Molinari (2020), and Kline and Tamer (2023) for detailed surveys of partial identification in economics. Relevant to this paper is partial identification with data combination, since when combining two different data sets the joint distribution between the data sets is often unobserved (Cross and Manski, 2002; Molinari and Peski, 2006; Ridder and Moffitt, 2007; Fan et al., 2014, 2016). However in the data combination literature the joint distribution within each data set is observed; in the aggregate data setting the joint distribution between every combination of variables is unobserved. In particular, the setting in Cross and Manski (2002) applies when we observe the marginal distribution of $Y_{i}$ over groups and the joint distribution of covariates $X_{i}=(X_{1i},\dots,X_{Li})$ over groups. Aggregate data often does not fit this setting; for example aggregate data can provide ethnicity distributions, gender distributions, education distributions, and income distributions but will not provide joint distributions of all of these variables, often due to privacy concerns. Being able to observe the joint distribution of covariates makes bounds narrower than when we do not observe the joint distribution, as further discussed in Appendix A.3. The intuition for this is that since individual-level effects usually depend on the joint distribution of all variables, being able to pin down the joint distribution of most of the variables will restrict the size of the identified set.
The rest of the paper proceeds as follows. Section 2 presents identified sets on the parameters of interest under a few different assumptions. Section 3 develops consistent estimation and inference procedures. Section 4 presents and discusses results from an empirical application using standardized exam data and simulation exercises calibrated to the dataset. Section 5 concludes.
2 Identification
2.1 Identified set
Let $(Y_{i},X_{1i},\dots,X_{Li},G_{i}),i=1,\dots,n$ be a sequence of random variables. Suppose outcome $Y_{i}$ is binary, with $Y_{i}\in\{0,1\}$ and suppose $G_{i}\in\{1,\dots,G\}$ denotes the group of individual $i$. Further assume covariates $X_{i}\equiv(X_{1i},\dots,X_{Li})^{\prime}$ are discrete with known finite support $\{x_{k}\}_{k=1}^{K}\subseteq\mathbb{R}^{L}$. [^5] The case with multinomial or continuous outcome and continuous covariates is beyond the scope of this paper.
The goal is to construct bounds on linear combinations of $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$, $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$, for given weights $\{\lambda_{k}\}_{k=1}^{K}$. For example, if we are interested in the average marginal effect of changing $X_{i}$ from $x_{k_{1}}$ to $x_{k_{2}}$ on $Y_{i}$, we can choose $\lambda_{k_{2}}=1,\lambda_{k_{1}}=-1,$ and $\lambda_{k}=0$ for all other $k$. I will construct identified sets using only expressions for $\mathbb{E}[Y_{i}|G_{i}=g],\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g],$ and $\mathbb{P}[G_{i}=g]$ for every $\ell=1,\dots,L,k=1,\dots,K,g=1,\dots,G$; the sample equivalents of these parameters are observed in aggregate data.
For example, in a data set of standardized exam results and demographics, $G_{i}$ denotes student $i$’s school district, $Y_{i}$ is an indicator for whether student $i$ passed the exam or not, and $X_{i}$ are student $i$’s demographics. We observe (sample estimates of) the pass rate for every school district $\mathbb{E}[Y_{i}|G_{i}]$, demographics of every school district $\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$, and the number of students enrolled in each school district, with which we can obtain (sample estimates of) $\mathbb{P}[G_{i}=g]$.
In what follows I will maintain the following assumption:
Assumption 1.
For random variables $Y_{i},G_{i}$, and $L$-dimensional random vector $X_{i}$, suppose
  1. 1.
    $Y_{i}$ is binary.
  2. 2.
    $G_{i}$ is discrete with finite support $\{1,\dots,G\}$.
  3. 3.
    $X_{i}$ is discrete with finite support $\{x_{k}\}_{k=1}^{K}\subseteq\mathbb{R}^{L}$.
  4. 4.
    $G_{i}$ is i.i.d. and $(Y_{i},X_{i})|G_{i}$ are i.i.d. [^6] I assume individual-level variables are (conditionally) i.i.d. for the sake of simplicity; this assumption can be relaxed and the consistency and inference results of Section 3 will still hold.
  5. 5.
    $(Y_{i},G_{i},X_{i})$ are latent; instead, we observe (sample analogs of) $\mathbb{E}[Y_{i}|G_{i}=g],\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g],$ and $\mathbb{P}[G_{i}=g]$ for every $\ell=1,\dots,L,k=1,\dots,K,g=1,\dots,G$.
Because $Y_{i}$ is binary, the law of total probability gives us
$\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ $=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\mathbb{E}[Y_{i}|X_{i}=x_{k}|G_{i}=g],$
$\mathbb{E}[Y_{i}|G_{i}=g]$ $=\sum_{k=1}^{K}\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]\mathbb{P}[X_{i}=x_{k}|G_{i}=g],$
$\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ $=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}\mathbb{P}[X_{i}=x_{j}|G_{i}=g].$
As in Cross and Manski (2002), these are the only relationships we can use to relate joint information of interest to the observed marginal information in the data without any further assumptions.
Let $\delta_{k}\equiv\mathbb{E}[Y_{i}|X_{i}=x_{k}]$, $\gamma_{kg}\equiv\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]$, and $\pi_{kg}\equiv\mathbb{P}[X_{i}=x_{k}|G_{i}=g]$ denote the unobserved parameters in the above equations. Then we can rewrite the equations as
$\delta_{k}$ $=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\gamma_{kg},$ (1)
$\mathbb{E}[Y_{i}|G_{i}=g]$ $=\sum_{k=1}^{K}\gamma_{kg}\pi_{kg}$ (2)
$\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ $=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}\pi_{jg}.$ (3)
Note that $\delta_{k},\gamma_{kg},\pi_{kg}\in[0,1]$ for all $k,g$. Combining this with equations (1), (2), and (3) above, we can define the identified set for $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$:
Lemma 1.
Suppose Assumption 1 holds. Given $\{\lambda_{k}\}_{k=1}^{K}\in\mathbb{R}^{K}$, the sharp identified set for $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ is given by
$D=\Bigg{\{}\sum_{k=1}^{K}\lambda_{k}d_{k}~{}\bigg{|}~{}0\leq d_{k}\leq 1~{}\forall k,\text{ and }\exists\,(p_{1g},\dots,p_{Kg})\in[0,1]^{K},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }d_{k}=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}~{}\forall k,~{}\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g,\\ \mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}~{}\forall\ell,k,g,\text{ and }\sum_{k=1}^{K}p_{kg}=1~{}\forall g\Bigg{\}}.$ (4)
Proposition 1.
$D=\left[\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]L_{g},\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]U_{g}\right]$, where
$L_{g}$ $\equiv\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},$
$U_{g}$ $\equiv\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},\text{ and}$
[…]
I defer all proofs to Appendix B. Proposition 1 states that we can equivalently express the identified set $D$ given by (4) as the weighted sum of solutions to bilevel optimization problems. This formulation is helpful because it suggests how computation of the lower and upper bound might be performed. In particular, solving for
$\min_{\{c_{kg}\}\in[0,1]^{K}}\underset{\{c_{kg}\}\in[0,1]^{K}}{(\max)}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ (5)
given a particular $\{p_{kg}\}\in P_{g}$ is a linear program. Thus we can solve for $L_{g}$ and $U_{g}$ by searching for the minimum and maximum respectively of the linear program (5) over all $\{p_{kg}\}\in P_{g}$. This optimization problem has a nonconvex objective; I suggest using existing derivative-free nonconvex solvers with a coarse grid of starting points to solve the optimization problem.
Remark 1.
I show in Appendix A.1 that the problem of solving the linear program (5) given any particular $\{p_{kg}\}\in P_{g}$ has an analytical solution. While computing the analytical solution for each given $\{p_{kg}\}$ is fast, computing the solution from the linear program formulation is also fast and either method can be used.
Remark 2.
If the parameter of interest is $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ instead of a linear combination, I demonstrate in Appendix A.2 that there exists an analytical solution to the sharp bounds using Fréchet inequalities.
Remark 3.
Without further restrictions on the underlying data generating process, it will always be the case that $\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|G_{i}=g]\in D$. To see why, note that, inspecting the optimization problems of Proposition 1, for any $g$ and any $\{p_{kg}\}\in P_{g}$, letting $c_{kg}=\mathbb{E}[Y_{i}|G_{i}=g]$ for all $k=1,\dots,K$ satisfies the constraint that $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ because $\sum_{k=1}^{K}p_{kg}=1$. This is relevant for average marginal effect parameters because the weights $\{\lambda_{k}\}$ are such that $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|G_{i}=g]=0$ for each $g$ so $0\in D$, and we thus cannot rule out a zero average marginal effect without further restrictions.
2.2 Additional shape restrictions
There may be situations in which the researcher is willing to assume polyhedral shape restrictions on the conditional expectation function over groups $\mathbb{E}[Y_{i}|X_{i},G_{i}]$. Examples of such shape restrictions include convexity, concavity, and monotonicity. [^7] See the reviews written by Matzkin (1994) and Chetverikov et al. (2018) for other examples of shape restrictions that have been used in econometric models. I impose the shape constraints as an additional assumption:
Assumption 2.
For each $g=1,\dots,G$,
$S_{g}Y_{X,g}\leq a_{g},$
where $Y_{X,g}\equiv\left(\mathbb{E}[Y_{i}|X_{i}=x_{1},G_{i}=g],\dots,\mathbb{E}[Y_{i}|X_{i}=x_{K},G_{i}=g]\right)^{\prime}$, $S_{g}\in\mathbb{R}^{s_{g}\times K}$ are known fixed matrices, and $a_{g}$ are known fixed vectors.
We can simply add this shape restriction to the constraints of the optimization problems solving $L_{g}$ and $U_{g}$ in Proposition 1, as in Freyberger and Horowitz (2015), to obtain sharp bounds on $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$:
Proposition 2.
Suppose Assumptions 1 and 2 hold. Given $\{\lambda_{k}\}_{k=1}^{K}\in\mathbb{R}^{K}$ and $S_{g}\in\mathbb{R}^{s_{g}\times K}$ for each $g$, the sharp identified set for $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ is given by
$D=\left[\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]L_{g},\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]U_{g}\right],$
where, defining $c_{g}\equiv(c_{1g},\dots,c_{Kg})^{\prime}$,
$L_{g}$ $\equiv\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }S_{g}c_{g}\leq a_{g}\text{ and }\exists\{p_{kg}\}\in P_{g}\text{with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},$
$U_{g}$ $\equiv\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }S_{g}c_{g}\leq a_{g}\text{ and }\exists\{p_{kg}\}\in P_{g}\text{with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},$
[…]
Again, given a particular $\{p_{kg}\}\in P_{g}$ the solution to
$\underset{\{c_{kg}\}\in[0,1]^{K}}{\min/\max}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }S_{g}c_{g}\leq a_{g}\text{ and }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ (6)
is a linear program. We can again solve for $L_{g}$ and $U_{g}$ by searching for the minimum and maximum respectively of (6), which is fast to compute, over all $\{p_{kg}\}\in P_{g}$. The suggested method of using a nonconvex solver with a coarse grid of starting points to solve the problem is still valid.
2.3 Additional aggregate data at a finer level
In some situations the researcher has access to additional data that is aggregated at a finer level than by groups $G_{i}$. In this section I consider the case when $\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$ is observed by the researcher. For example, in a data set of standardized exam results and demographics, the researcher may have access to average pass results by race in some school districts.
Assumption 3.
We observe (sample analogs of) $\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$ for all $(\ell,k)\in F_{g}$, where for each group $G_{i}=g$, $F_{g}$ is a (possibly empty) set of indices $(\ell,k)$.
By the definition of conditional probability and the law of total probability, for each $(\ell,k)\in F_{g}$,
$\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$
$=\mathbb{P}[Y_{i}=1,X_{\ell i}=x_{k,\ell}|G_{i}=g]$
$=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]$
$=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]\mathbb{P}[X_{i}=x_{k}|G_{i}=g]$
$=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}\gamma_{kg}\pi_{kg}.$
Again, we can easily add these restrictions to the constraints of the optimization problem solving $L_{g}$ and $U_{g}$ in Proposition 1 to obtain sharp bounds on $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$:
Proposition 3.
Suppose Assumptions 1 and 3 hold. Given $\{\lambda_{k}\}_{k=1}^{K}\in\mathbb{R}^{K}$, the sharp identified set for $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ is given by
$D=\left[\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]L_{g},\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]U_{g}\right],$
where
$L_{g}$ $\equiv\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}\text{ and}$
$\qquad\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g},$
$U_{g}$ $\equiv\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}\text{ and}$
$\qquad\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g},$
[…]
Since the additional constraints are all linear constraints, the solution to $L_{g}$ and $U_{g}$ for a particular $\{p_{kg}\}\in P_{g}$ is still a linear program. The previous discussion about how to compute the identified set applies here.
Remark 4.
The sharp identified set $D$ under Assumptions 1, 2, and 3 is given by adding both of the restrictions $S_{g}c_{g}\leq a_{g}$ and
$\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g}$
to the optimization problems of $L_{g}$ and $U_{g}$ in Proposition 1.
Remark 5.
This methodology is flexible and can easily be adjusted to accommodate further assumptions beyond Assumptions 2 and 3 on the underlying individual-level model through additional restrictions to the optimization problem. For example, if some of the covariances between covariates can be estimated or bounded using a separate data set, this could be incorporated as polyhedral restrictions on the joint support of the covariates in the $P_{g}$ optimization problem. Restrictions on the underlying distribution of $X_{i}$ can also be incorporated through specification of the support.
Remark 6.
If we impose the assumption that the $Y_{i},X_{1i},\dots,X_{Li}$ are mutually independent conditional on $G_{i}$, we would obtain point identification of $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ and thus point identification of $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$. If we impose that the $X_{1i},\dots,X_{Li}$ are mutually independent conditional on $G_{i}$, we would obtain point identification of the joint distribution of $X_{i}|G_{i}$. Sharp bounds on $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ would then follow from Cross and Manski (2002): we can consider the $L_{g}$ and $U_{g}$ optimization problems but letting $p_{kg}$ be the joint distribution of $X_{i}|G_{i}$.
3 Estimation and Inference
3.1 Estimation
In practice we observe sample analogs of the population values $\mathbb{E}[Y_{i}|G_{i}=g],\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g],\mathbb{P}[G_{i}=g]$ in the aggregate data set, along with sample analogs of $\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$ if available. For all $\ell=1,\dots,L,j=1,\dots,K,g=1,\dots,G$, denote
$\bar{Y}_{g}$ $=\frac{\sum_{i=1}^{n}Y_{i}\mathbbm{1}\{G_{i}=g\}}{\sum_{i=1}^{n}\mathbbm{1}\{G_{i}=g\}}$ (7)
$\widehat{Pr}[X_{\ell i}=x_{j,\ell}|G_{i}=g]$ $=\frac{\sum_{i=1}^{n}\mathbbm{1}\{X_{\ell i}=x_{j,\ell}\}\mathbbm{1}\{G_{i}=g\}}{\sum_{i=1}^{n}\mathbbm{1}\{G_{i}=g\}}$ (8)
$\widehat{Pr}[G_{i}=g]$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{G_{i}=g\}$ (9)
$\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}$ $=\frac{\sum_{i=1}^{n}Y_{i}\mathbbm{1}\{X_{\ell i}=x_{k,\ell}\}\mathbbm{1}\{G_{i}=g\}}{\sum_{i=1}^{n}\mathbbm{1}\{X_{\ell i}=x_{k,\ell}\}\mathbbm{1}\{G_{i}=g\}}.$ (10)
$\bar{Y}_{g},\widehat{Pr}[X_{\ell i}=x_{j,\ell}],\widehat{Pr}[G_{i}=g],$ and $\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}$ all converge to the respective population values by the law of large numbers and continuous mapping theorem, assuming group probabilities $\mathbb{P}[G_{i}=g]$ and covariate probabilities $\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ are bounded away from 0. Thus we can construct a plug-in estimator, denote $\hat{D}$, for the sharp identified set by replacing all population values in the optimization problems of Propositions 1, 2, 3, or Remark 4 with their sample estimates. For example, the estimated sharp identified set discussed in Remark 4 looks like:
$\hat{D}\equiv\left[\hat{L},\hat{U}\right]\equiv\left[\sum_{g=1}^{G}\widehat{Pr}[G_{i}=g]\hat{L}_{g},\sum_{g=1}^{G}\widehat{Pr}[G_{i}=g]\hat{U}_{g}\right],$
$\hat{L}_{g}$ $\equiv\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }S_{g}c_{g}\leq a_{g}~{}\forall g\text{ and }\exists\{p_{kg}\}\in\hat{P}_{g}\text{ with }\bar{Y}_{g}=\sum_{k=1}^{K}c_{kg}p_{kg},$
$\qquad\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}\times\widehat{Pr}[X_{\ell i}=x_{j,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g},$
$\hat{U}_{g}$ $\equiv\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }S_{g}c_{g}\leq a_{g}~{}\forall g\text{ and }\exists\{p_{kg}\}\in\hat{P}_{g}\text{ with }\bar{Y}_{g}=\sum_{k=1}^{K}c_{kg}p_{kg},$
$\qquad\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}\times\widehat{Pr}[X_{\ell i}=x_{j,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g},$
$\hat{P}_{g}\equiv\text{arg}\min_{\{p_{kg}\}}\sum_{r=1}^{LK}v_{r}^{+}+v_{r}^{-}\;\;\text{s.t.}\sum_{k=1}^{K}p_{kg}=1;p_{kg},v_{r}^{+},v_{r}^{-}\geq 0~{}\forall k,r;\text{ and }\\[-10.0pt] \widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]-\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}=v_{K(\ell-1)+k}^{+}-v_{K(\ell-1)+k}^{-}~{}\forall\ell,k.$
The following proposition shows that the lower and upper bounds of the plug-in estimated set $\hat{D}$ are consistent.
Proposition 4.
Suppose Assumption 1 holds and that $\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ are valid marginal distributions for $X_{i}$ with respect to the assumed support. If Assumption 2 holds, define $\hat{D}$ with respect to Proposition 2; if Assumption 3 holds, define $\hat{D}$ with respect to Proposition 3; and if Assumptions 2 and 3 hold, define $\hat{D}$ with respect to Remark 4.
[…]
3.2 Inference
Existing inference methods for partially identified estimated sets usually require knowledge of the joint distribution of the individual-level data to estimate a covariance matrix used in constructing critical values or test statistics for valid coverage. However, in my setting I only observe marginal distributions of each variable. I point out why existing methods cannot be applied for a few examples, by no means representative, below:
Example.
Horowitz and Manski (2000) derive analytic bounds on conditional mean outcomes and point out that the delta method delivers asymptotic normality of the lower and upper bound estimators. The paper bootstraps the asymptotic covariance matrix to obtain a confidence interval that contains the identified set with correct asymptotic coverage. Putting aside that I do not have analytic bounds in my setting, being able to derive the asymptotic covariance of bounds $\hat{L}_{g}$ and $\hat{U}_{g}$ requires that I know the covariances between, for example, $Y_{i}$ and any $X_{\ell i}$ or any $X_{{\ell_{1}}i}$ and $X_{{\ell_{2}}i}$ given $G_{i}$. However I only observe sample marginal distributions of each variable and thus cannot hope to estimate the covariance matrix. Bootstrapping will also not be possible because I do not observe the individual-level data and so cannot generate a bootstrap sample that reflects the dependence between all of the variables.
Example.
Imbens and Manski (2004) provide a method for constructing confidence intervals on the parameter value of interest instead of on the entire identified set. This method chooses critical values for correct coverage by again relying on joint asymptotic normality of the lower and upper bound estimators, $\hat{L}_{g}$ and $\hat{U}_{g}$ in my setting. As discussed in the above example, I cannot hope to estimate the variances of the lower and upper bounds with the marginal information observed in aggregate data alone.
Example.
Hsieh et al. (2022) construct confidence intervals for identified sets of solutions to convex optimization programs, specifically linear and quadratic optimization programs with estimated coefficients, exploiting the necessary and sufficient optimality conditions. Putting aside that the optimization program is nonconvex in my setting, implementation of this inference method requires the asymptotic covariance matrix of the estimated covariates of the optimization problem, which in my setting rely on $\bar{Y}_{g},\widehat{Pr}[X_{\ell i}=x_{j,\ell}|G_{i}=g],$ and $\widehat{Pr}[G_{i}=g]$. Thus I again need to know the covariances between, for example, $Y_{i}$ and any $X_{\ell i}$ or any $X_{{\ell_{1}}i}$ and $X_{{\ell_{2}}i}$ given $G_{i}$.
Therefore in this setting I am forced to rely on inference methods that require only marginal information of each variable. I choose to use the Bonferroni correction to make marginal confidence intervals on each sample observation jointly valid across the whole sample. Since each aggregate observation is the sample average of a binary random variable, as can be seen from equations (7), (8), (9), and (10), I can use Clopper-Pearson intervals, which are finite-sample valid, for each sample observation instead of relying on normal approximations. [^8] Note that any other binomial proportion confidence interval that obtains asymptotically nominal coverage will provide asymptotically valid coverage; Clopper-Pearson has the advantage of being finite-sample valid.
Let $M$ be the total number of observations in the aggregate data set, that is, the total number of $\bar{Y}_{g},\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g],\widehat{Pr}[G_{i}=g]$, and $\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}$ (if observed) observations in the data sets across all groups $g$, support points $k$ and covariates $\ell$.
The inference procedure is as follows:
  1. 1.
    For every sample observation $\hat{p}$, construct two-sided level $1-\frac{\alpha}{M}$ Clopper-Pearson CIs, denoted $\left[\hat{p}_{L},\hat{p}_{U}\right]$. For $N$ the number of individuals in the conditioning set of the population analog of $\hat{p}$ (i.e. for $\bar{Y}_{g}$, $N=\sum_{i=1}^{n}\mathbbm{1}\{G_{i}=g\}$ and for $\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}$, $N=\sum_{i=1}^{n}\mathbbm{1}\{X_{\ell i}=x_{k,\ell}\}\mathbbm{1}\{G_{i}=g\}$) and $X=N\hat{p}$, the Clopper-Pearson CI is determined by quantiles of the beta distribution:
    $\left[\hat{p}_{L},\hat{p}_{U}\right]=\left[B\bigg{(}\frac{\alpha}{2M},X,N-X+1\bigg{)},B\bigg{(}1-\frac{\alpha}{2M},X+1,N-X\bigg{)}\right]$
    The resulting confidence intervals are $\left[\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]_{L},\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]_{U}\right]$ for each $\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$, $\left[\bar{Y}_{g,L},\bar{Y}_{g,U}\right]$ for each $\bar{Y}_{g}$, $\left[\widehat{Pr}[G_{i}=g]_{L},\widehat{Pr}[G_{i}=g]_{U}\right]$ for each $\widehat{Pr}[G_{i}=g]$, and (if observed) $\left[{\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}}_{L},{\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}}_{U}\right]$ for each $\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}$.
  2. 2.
    Solve the optimization programs of $\hat{D}$ under the assumed assumptions for all values of sample observations within the marginal confidence intervals constructed in step 1. For example, under both Assumptions 1, 2, and 3, we solve:
    $a)\quad\hat{P}_{g,CI}\equiv\text{arg}\min_{\{p_{kg}\}}\sum_{r=1}^{2LK}v_{r}^{+}+v_{r}^{-}\;\;\text{s.t.}\sum_{k=1}^{K}p_{kg}=$  $\;1;p_{kg},v_{r}^{+},v_{r}^{-}\geq 0~{}\forall k,r;$
    $\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]_{L}-\sum_{k=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}\leq$ $\;v_{K(\ell-1)+k}^{+}-v_{K(\ell-1)+k}^{-}~{}\forall\ell,k,\text{ and }$
    $\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]_{U}-\sum_{k=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}$ $\geq v_{K(L+\ell-1)+k}^{+}-v_{K(L+\ell-1)+k}^{-}~{}\forall\ell,k.$
    $b)\quad\hat{L}_{g,CI}\equiv\min_{\{c_{gk}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{gk}\text{ s.t. }S_{g}c_{g}\leq a_{g}~{}\forall g\text{ and }\exists\{p_{gk}\}\in\hat{P}_{g,CI}\\[-10.0pt] \text{with }\bar{Y}_{g,L}\leq\sum_{k=1}^{K}c_{gk}p_{gk},\bar{Y}_{g,U}\geq\sum_{k=1}^{K}c_{gk}p_{gk},\\[-8.0pt] {\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}}_{L}\times\widehat{Pr}[X_{\ell i}=x_{j,\ell}|G_{i}=g]_{L}\leq\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g},\\[-11.0pt] \text{ and }{\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}}_{U}\times\widehat{Pr}[X_{\ell i}=x_{j,\ell}|G_{i}=g]_{U}\geq\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g}$
    $c)\quad\hat{U}_{g,CI}\equiv\max_{\{c_{gk}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{gk}\text{ s.t. }S_{g}c_{g}\leq 0~{}\forall g\text{ and }\exists\{p_{gk}\}\in\hat{P}_{g,CI}\\[-10.0pt] \text{with }\bar{Y}_{g,L}\leq\sum_{k=1}^{K}c_{gk}p_{gk},\bar{Y}_{g,U}\geq\sum_{k=1}^{K}c_{gk}p_{gk},\\[-8.0pt] {\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}}_{L}\times\widehat{Pr}[X_{\ell i}=x_{j,\ell}|G_{i}=g]_{L}\leq\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g},\\[-11.0pt] \text{ and }{\bar{Y}_{g}|X_{\ell i}=x_{k,\ell}}_{U}\times\widehat{Pr}[X_{\ell i}=x_{j,\ell}|G_{i}=g]_{U}\geq\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g}.\vspace{0pt}$
  3. 3.
    The confidence interval is given by
    $\hat{D}_{CI}\equiv\left[\sum_{g=1}^{G}\widehat{Pr}[G_{i}=g]_{L}\hat{L}_{g,CI},\sum_{g=1}^{G}\widehat{Pr}[G_{i}=g]_{U}\hat{U}_{g,CI}\right].$
Proposition 5.
Suppose Assumption 1 holds. If Assumption 2 holds, define $\hat{D}_{CI}$ with respect to Proposition 2; if Assumption 3 holds, define $\hat{D}_{CI}$ with respect to Proposition 3; and if Assumptions 2 and 3 hold, define $\hat{D}_{CI}$ with respect to Remark 4.
Suppose also that the $\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ are valid marginal distributions for $X_{i}$ with respect to the assumed support. Then $\mathbb{P}[D\subseteq\hat{D}_{CI}]\geq 1-\alpha$.
Since $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]\in D$, this means that $\mathbb{P}[\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[Y_{i}|X_{i}=x_{k}]\in\hat{D}_{CI}]\geq 1-\alpha$, as discussed in Imbens and Manski (2004). Proposition 5 says that the confidence interval $\hat{D}_{CI}$ has correct coverage for the identified set and thus for the identified parameter.
4 Simulations and Empirical Application
One setting in which publicly available data is in aggregate form is standardized exam data. In this section I will apply the methodology developed in the previous sections to construct bounds on conditional exam pass rates. In this application I focus on exam pass rates for English and math RICAS exams and student demographic information for the state of Rhode Island in spring of 2018 and spring of 2019 over all students in grades 3-8.999Much of this data is a subset of data from Jack et al. (2023) in their paper studying the effect of online learning during the COVID-19 pandemic on exam results. I thank the authors for their assistance with obtaining the data.
In my application I produce sharp bounds on the average pass rate conditional on three covariates: race (white versus non-white), eligibility for free and reduced price lunch (FRPL), and English-language learner (ELL) status. In Section 4.1 I first explore what causes the width of the bounds to vary in simulations calibrated to the Rhode Island aggregate data. In Section 4.2 I then present the empirical application, where I estimate bounds with and without monotonicity shape restrictions and additional pass rate data for subgroups.
4.1 Simulations
I present three different simulation exercises. In all exercises I have three binary covariates $white_{i},FRPL_{i}$, and $ELL_{i}$, and a binary outcome $pass_{i}$. There are 58 aggregate groups in each example, where in the Rhode Island data a group is a school district-year pair (each school district in a given year), and in all simulation exercises I assume there are 2000 individuals in each group.
In the first exercise, I choose a joint distribution such that the marginal distribution over groups of each covariate approximately matches the marginal distribution over groups of each aggregate-level covariate in the Rhode Island data, as plotted in Figure 1.
Figure 1: Distributions of aggregate variables in simulation exercise 1
I present results for 100 different aggregate data sets generated according to the joint distribution of exercise 1 in Figure 2. The parameters for which I produce bounds in this figure are average conditional outcomes $\mathbb{E}[pass_{i}|white_{i},FRPL_{i},ELL_{i}]$. The 95% confidence intervals contain the population bounds for all 100 data sets and all parameters, suggesting that the confidence intervals are conservative, as to be expected with the Bonferroni correction and Clopper-Pearson intervals. The estimated bounds are very close to the population bounds for all parameters. Because of this observation, I will sometimes focus on results for one particular simulated data set and only present estimated bounds without corresponding confidence intervals.
Figure 2: Estimated bounds and 95% CIs over 100 draws in simulation exercise 1
Table 1 presents results from the first simulation exercise for one of the aggregate data sets generated according to the joint distribution of exercise 1. Population bounds in Column 2 are wide, with most of them uninformative (equal to $[0,1]$). Estimated bounds in Column 3 are very close to the population bounds in Column 2. The 95% confidence intervals in Column 4 are a bit wider than the estimated bounds. The parameters for which I obtain informative bounds seem to be those where the conditioning population is well-represented in the data: the simulated data is mostly groups with a large fraction of white students and small fractions of FRPL and ELL students, and the parameter with the tightest bounds is the average pass rate among white, non-FRPL, non-ELL students. Sharp bounds on the difference between parameters are the Minkowski set difference between bounds on each of the parameters. In particular, as noted in Remark 3, all bounds on the difference between parameters contain 0 as I do not impose additional assumptions.
Table 1: Estimated bounds on conditional pass rate in simulation exercise 1
Population Estimated Estimated Bounds
True Value Bounds Bounds 95% CI on Difference
Parameter (1) (2) (3) (4) (5)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $0.618$ [$0.179,0.860$] $[0.180,0.860]$ $[0.126,0.954]$ $[-0.820,0.860]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$ $0.5$ $[0,1]$ $[0,1]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $0.274$ [$0,0.980$] $[0,0.981]$ $[0,0.986]$ $[-0.997,0.981]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $0.184$ [$0,0.990$] $[0,0.997]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $0.460$ $[0,1]$ $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $0.345$ $[0,1]$ $[0,1]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $0.159$ $[0,1]$ $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $0.097$ $[0,1]$ $[0,1]$ $[0,1]$
Notes: 95% CIs are confidence intervals on the estimated bounds in Column 3. CIs are constructed using the method from Section 3.2, taking $\mathbb{P}[G_{i}=g]$ as observed instead of $\widehat{Pr}[G_{i}=g]$. Bounds on the difference are sharp bounds on the top minus the bottom parameter for which each set of bounds are reported.
To see if the estimated bounds could be narrower, in the second simulation exercise I choose a joint distribution that produces the same true conditional average pass rates, but such that the marginal distribution over groups of each covariate is even closer to either 0 or 1, as can be seen in Figure 3. Results for 100 different aggregate data sets generated according to the joint distribution are in Figure 4. Again, the 95% confidence intervals contain the population bounds for all 100 data sets and the estimated bounds are very close to the population bounds.
Figure 3: Distributions of aggregate variables in simulation exercise 2
Figure 4: Estimated bounds and 95% CIs over 100 draws in simulation exercise 2
Results for one aggregate data set generated according to the joint distribution of exercise 2 are presented in Table 2. In Column 3, bounds on the parameter for which the conditioning population is well-represented in the data are narrower and more informative than in the first exercise, but bounds on all other parameters are uninformative, while in the first exercise some of the bounds on other parameters were informative. However, bounds on the difference in the first row of Column 5 are still wide, even though they are narrower than in the first exercise. This is likely because while there is more information about the first parameter, there is less information on the second parameter so that bounds on the difference are still wide.
Table 2: Estimated bounds on conditional pass rate in simulation exercise 2
Population Estimated Bounds on
True Value Bounds Bounds 95% CI Difference
Parameter (1) (2) (3) (4) (5)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $0.618$ [$0.572,0.749$] $[0.572,0.749]$ $[0.501,0.843]$ $[-0.428,0.749]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$ $0.5$ $[0,1]$ $[0,1]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $0.274$ $[0,1]$ $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $0.184$ $[0,1]$ $[0,1]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $0.460$ $[0,1]$ $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $0.345$ $[0,1]$ $[0,1]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $0.159$ $[0,1]$ $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $0.097$ $[0,1]$ $[0,1]$ $[0,1]$
Notes: 95% CIs are confidence intervals on the estimated bounds in Column 3. CIs are constructed using the method from Section 3.2, taking $\mathbb{P}[G_{i}=g]$ as observed instead of $\widehat{Pr}[G_{i}=g]$. Bounds on the difference are sharp bounds on the top minus the bottom parameter for which each set of bounds are reported.
This suggests that if the conditioning populations for the first two parameters were well-represented, bounds on the difference might be narrower. In the third simulation exercise I choose a joint distribution that produces the same true conditional average pass rates and marginal distributions over groups for $\mathbb{E}[FRPL_{i}]$ and $\mathbb{E}[ELL_{i}]$, but I let some groups have $\mathbb{E}[white_{i}]$ close to 0 and others have $\mathbb{E}[white_{i}]$ close to 1, as can be seen in Figure 5. Results for 100 different aggregate data sets generated according to this joint distribution are in Figure 6. As in the previous two exercises, the 95% confidence intervals contain the population bounds for all 100 data sets and the estimated bounds are very close to the population bounds.
Figure 5: Distributions of aggregate variables in simulation exercise 3
Figure 6: Estimated bounds and 95% CIs over 100 draws in simulation exercise 3
Results for one aggregate data set generated according to the joint distribution of exercise 3 are presented in Table 3. As expected, bounds on each parameter in Column 3 are relatively narrow on the parameters for which the conditioning population is well-represented in the data. Notably relative to the second exercise, bounds on the first parameter are wider and bounds on the second parameter are narrower. This suggests that there is a trade-off between obtaining narrow bounds on a single parameter and obtaining narrow bounds on multiple parameters. Intuitively, the groups with $\mathbb{E}[white_{i}]$ close to 0 help to make bounds on the second parameter narrow but make bounds on the first parameter wider.
Table 3: Estimated bounds on conditional pass rate in simulation exercise 3
Population Estimated Bounds on
True Value Bounds Bounds 95% CI Difference
Parameter (1) (2) (3) (4) (5)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $0.618$ [$0.315,0.860$] $[0.316,0.860]$ $[0.277,0.913]$ $[-0.222,0.860]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$ $0.5$ [$0,0.539$] $[0,0.539]$ $[0,0.555]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $0.274$ $[0,1]$ $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $0.184$ $[0,1]$ $[0,1]$ $[0,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $0.460$ $[0,1]$ $[0,1]$ $[0,1]$ $[-0.897,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $0.345$ [$0,0.907$] $[0,0.897]$ $[0,0.949]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $0.159$ $[0,1]$ $[0,1]$ $[0,1]$ $[-1,1]$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $0.097$ $[0,1]$ $[0,1]$ $[0,1]$
Notes: 95% CIs are confidence intervals on the estimated bounds in Column 3. CIs are constructed using the method from Section 3.2, taking $\mathbb{P}[G_{i}=g]$ as observed instead of $\widehat{Pr}[G_{i}=g]$. Bounds on the difference are sharp bounds on the top minus the bottom parameter for which each set of bounds are reported.
The bounds on the difference in the first row of Column 5 are narrower than in the first or the second exercise, but are still wide and, important to signing the parameter, contain 0. This means that, from this simulated data set, it would not be possible to say whether the white/non-white average pass rate gap among any group of FRPL/non-FRPL and ELL/non-ELL students is positive or negative. This suggests that without further restrictions, obtaining usefully informative bounds on average marginal effects is challenging.
4.2 Empirical application
The results of the simulation exercise suggest that bounds on white/non-white average pass rate gaps will not be informative in the Rhode Island data. I thus consider imposing several additional assumptions to see whether more restrictions and information can help make bounds narrower.
4.2.1 Monotonicity restrictions
I first consider imposing several monotonicity shape restrictions. Motivated by test score gaps that have been documented between rich and poor students (Tavernise, 2012; Porter, 2015), I impose that for each value of $(white_{i},ELL_{i})$ and each the average pass rate is lower for FRPL students than for non-FRPL students:
$\mathbb{E}[pass_{i}|FRPL_{i}=1,white_{i},ELL_{i},G_{i}=g]-\mathbb{E}[pass_{i}|FRPL_{i}=0,white_{i},ELL_{i},G_{i}=g]\leq 0.$ (11)
I first present estimated bounds on math exam white/non-white average pass rate gaps in Table 4. As expected, sharp bounds on the white/non-white average pass rate gaps reported in Column 1 are wide and either uninformative or close to uninformative. Imposing the additional monotonicity restriction (11) helps narrow the bounds for some parameters, reported in Column 2, but bounds are still wide and contain 0.
Table 4: Rhode Island white/non-white math exam pass rate differences
Bounds without Monotonicity Monotonicity Bounds
Parameter (1) (2)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.940,0.849]$ $[-0.873,0.849]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-0.816,0.994]$ $[-0.687,0.599]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-1,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-0.912,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$
Notes: Bounds on difference impose no shape restrictions; monotonicity bounds impose monotonicity restriction (11), as discussed in Section 4.2.
For English exams, I impose an additional monotonicity restriction that for each value of $(white_{i},FRPL_{i})$ the average pass rate is lower for English-language learner students than for non-English-language learner students:
$\mathbb{E}[pass_{i}|ELL_{i}=1,white_{i},FRPL_{i},G_{i}=g]-\mathbb{E}[pass_{i}|ELL_{i}=0,white_{i},FRPL_{i},G_{i}=g]\leq 0.$ (12)
I present estimated bounds on English exam white/non-white average pass rate gaps in Table 5. Sharp bounds on the white/non-white average pass rate gaps reported in Column 1 are again wide and either uninformative or close to uninformative. Imposing both additional monotonicity restrictions of (11) and (12) helps to narrow the bounds on all parameters, reported in Column 2, but all bounds are still wide and contain 0.
Table 5: Rhode Island white/non-white English exam pass rate differences
Bounds without Monotonicity Monotonicity Bounds
Parameter (1) (2)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.903,0.923]$ $[-0.816,0.909]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-0.849,1]$ $[-0.755,0.708]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-1,0.913]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-0.716,0.657]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$
Notes: Bounds on difference impose no shape restrictions; monotonicity bounds impose monotonicity restrictions (11) and (12), as discussed in Section 4.2.
4.2.2 Additional pass rates among subgroups
Rhode Island makes available RICAS assessment results aggregated by student subgroup publicly available on their data portal. In particular, exam pass rates at the district level for white/nonwhite students are available for all school districts.
I present bounds using the additional subgroup data on the math exam white/non-white average pass rate gaps in Table 6. Bounds using the additional subgroup data on the English exam white/non-white average pass rate gaps are presented in Table 7. I present bounds using the additional data both without and with the monotonicity assumptions of Section 4.2.1.
Table 6: Rhode Island white/non-white math exam pass rate differences, with subgroup data
Bounds without Monotonicity Monotonicity Bounds
Parameter (1) (2)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.848,0.745]$ $[-0.765,0.641]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-0.740,0.968]$ $[-0.269,0.499]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-1,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-0.892,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$
Notes: Bounds on difference impose no shape restrictions but use additional pass rate by subgroup data; monotonicity bounds impose monotonicity restriction (11) in additional to additional pass rate by subgroup data. See Section 4.2 for more details.
Without the monotonicity assumptions, using the additional subgroup data makes the estimated bounds slightly narrower than without the additional subgroup data, as we can see from comparing Column 1 of Tables 4 and 6 and comparing Column 1 of Tables 5 and 7. Bounds under the monotonicity assumptions are also narrower with the additional subgroup data than without, as we can see from comparing Column 2 of Tables 4 and 6 and comparing Column 2 of Tables 5 and 7. The additional monotonicity assumption for English exam data appears to make bounds significantly narrower, especially in combination with the additional subgroup data. Although all bounds still contain 0, these results seem to suggest that the value of having additional information in the form of finer levels of aggregation goes some way towards obtaining more information about marginal effect parameters of interest, especially in combination with other shape restrictions.
Table 7: Rhode Island white/non-white English exam pass rate differences, with subgroup data
Bounds without Monotonicity Monotonicity Bounds
Parameter (1) (2)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.796,0.805]$ $[-0.590,0.567]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-0.767,0.981]$ $[-0.349,0.575]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-0.996,0.856]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-0.303,0.408]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$
Notes: Bounds on difference impose no shape restrictions but use additional pass rate by subgroup data; monotonicity bounds impose monotonicity restrictions (11) and (12) in additional to additional pass rate by subgroup data. See Section 4.2 for more details.
5 Conclusion
In this paper I present sharp bounds on individual-level parameters of interest when the only available data is aggregate data, and I develop a valid inference method relying only on the available marginal information of each covariate. In simulations and an empirical application I show that the sharp bounds are too wide to be useful with realistic aggregate data. Both additional shape restrictions and using additional data at a finer level of aggregation help to make sharp bounds narrower and together significantly narrow the sharp bounds. However, sharp bounds are unable to pin down, for example, the signs of marginal effects. This suggests that individual-level analyses using aggregate data are more precise when the aggregate data is available at a finer level of aggregation and when there is more underlying structure known about the individual-level data generating process. Researchers hoping to use aggregate data to make individual-level claims should try to obtain additional data at as fine of a level as possible and impose carefully justified assumptions in order to obtain useful results.
Appendix A Appendix: Additional Results
A.1 Closed-form solution for $L_{g}$ and $U_{g}$ problems
Not only is (5) a linear program, the following corollary shows (5) has a closed-form solution given any $\{p_{kg}\}\in P_{g}$:
Corollary A.1.
For given weights $\lambda_{1},\dots,\lambda_{K}$ and any fixed $\{p_{kg}\}\in P_{g}$, relabel the indices $k=1,\dots,K$ so that $\frac{\lambda_{1}}{p_{1g}}\geq\dots\geq\frac{\lambda_{K}}{p_{Kg}}$, where if $p_{kg}=0$ we define $\frac{\lambda_{k}}{p_{gk}}\equiv+\infty$. Then
$\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$
is attained by letting
$c_{1g}$ =$=\begin{cases}\max\left\{0,\frac{\mathbb{E}[Y_{i}|G_{i}=g]-1+p_{1g}}{p_{1g}}\right\}&p_{1g}>0\\ 0&p_{1g}=0\end{cases},$
$c_{kg}$ =$=\begin{cases}0&\sum_{j=1}^{k}p_{jg}\leq 1-\mathbb{E}[Y_{i}|G_{i}=g]\\ \frac{\mathbb{E}[Y_{i}|G_{i}=g]-1+\sum_{j=1}^{k}p_{jg}}{p_{kg}}&\sum_{j=1}^{k-1}p_{jg}\leq 1-\mathbb{E}[Y_{i}|G_{i}=g]<\sum_{j=1}^{k}p_{jg}\\ 1&\sum_{j=1}^{k-1}p_{jg}>1-\mathbb{E}[Y_{i}|G_{i}=g]\end{cases},$
and
$\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$
is attained by letting
$c_{1g}$ =$=\begin{cases}\min\left\{1,\frac{\mathbb{E}[Y_{i}|G_{i}=g]}{p_{1g}}\right\}&p_{1g}>0\\ 1&p_{1g}=0\end{cases}$
$c_{kg}$ =$=\begin{cases}1&\sum_{j=1}^{k}p_{jg}\leq\mathbb{E}[Y_{i}|G_{i}=g]\\ \frac{\mathbb{E}[Y_{i}|G_{i}=g]-\sum_{j=1}^{k-1}p_{jg}}{p_{kg}}&\sum_{j=1}^{k-1}p_{jg}\leq\mathbb{E}[Y_{i}|G_{i}=g]<\sum_{j=1}^{k}p_{jg}\\ 0&\sum_{j=1}^{k-1}p_{jg}>\mathbb{E}[Y_{i}|G_{i}=g]\end{cases}.$
Proofs are collected in Appendix B.
A.2 Fréchet inequalities
In Section 2.1 I derived sharp bounds for a linear combination of $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$, defined using the solutions to optimization programs. In this section I show that if we are interested in obtaining sharp bounds on each $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$ parameter under Assumption 1 alone, we can obtain closed-form bounds using Fréchet inequalities.
The Fréchet inequalities, explicitly derived by Fréchet (1951), state that if there are $N$ events $A_{1},\dots,A_{N}$, it holds that
$\max\left\{1-N+\sum_{i=1}^{N}Pr[A_{i}],0\right\}\leq Pr\left[\bigcap_{i=1}^{N}A_{i}\right]\leq\min\left\{Pr[A_{1}],\dots,Pr[A_{N}]\right\}.$ (A.1)
Therefore, for $y$ in the support of $Y_{i}$ and $x_{k}$ in the support of $X_{i}$, it follows from the Fréchet inequalities that
$L_{g}(y,x_{k})$ $\equiv\max\left\{\mathbb{P}[Y_{i}=y|G_{i}=g]+\sum_{\ell=1}^{L}\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]-L,0\right\}$
$\leq\mathbb{P}\left[Y_{i}=y,X_{i}=x_{k}|G_{i}=g\right]$
$\leq\min\left\{\mathbb{P}[Y_{i}=y|G_{i}=g],\mathbb{P}[X_{1i}=x_{k,1}|G_{i}=g],\dots,\mathbb{P}[X_{Li}=x_{k,L}|G_{i}=g]\right\}$
$\equiv U_{g}(y,x_{k}).$ (A.2)
Note that we only require knowledge of the marginal probabilities of each random variable in order to be able to calculate the lower and upper bounds; joint probabilities are not needed. If we are only given $\mathbb{P}[Y_{i}=y|G_{i}=g],\mathbb{P}[X_{1i}=x_{1}|G_{i}=g],\dots,\mathbb{P}[X_{Li}=x_{L}|G_{i}=g]$ and nothing else, then it is well known that the Fréchet inequalities are sharp; that is, they are the tightest possible bounds given the assumptions. Situations in which the Fréchet inequalities are not sharp include those in which we know variables are independent or if the (known) support of the variables is such that knowledge of a marginal probability provides knowledge of the joint probability.
For a given $k$ we can define bounds on $\mathbb{E}[Y_{i}|X_{i}=x_{k}]$:
Proposition A.1.
Suppose Assumption 1 holds. Let
$D^{F}\equiv\left[\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]L^{F}_{g},\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]U^{F}_{g}\right],$
where
$L^{F}_{g}$ $\equiv\sum_{k=1}^{K}\lambda_{k}\frac{L_{g}(1,x_{k})}{L_{g}(1,x_{k})+U_{g}(0,x_{k})},$
$U^{F}_{g}$ $\equiv\sum_{k=1}^{K}\lambda_{k}\frac{U_{g}(1,x_{k})}{L_{g}(0,x_{k})+U_{g}(1,x_{k})}.$
Then $D\subseteq D^{F}$, and if the Fréchet inequalities on $\mathbb{P}[Y_{i}=y,X_{i}=x_{k}|G_{i}=g]$ are sharp for all $y\in\{0,1\},k=1,\dots,K,g=1,\dots,G$ then $D=D^{F}$.
To construct an estimator of the identified set $D^{F}$, the sample analogs of $\mathbb{P}[Y_{i}=y|G_{i}=g],\mathbb{P}[X_{1i}=x_{1}|G_{i}=g],\dots,\mathbb{P}[X_{Li}=x_{L}|G_{i}=g]$ observed in the aggregate data can be plugged into the formula given in Proposition A.1. These estimated lower and upper bounds will be consistent because they are numerically equivalent to the estimated bounds of Proposition 1, which are consistent by Proposition 4.
To construct a valid confidence region for the identified set $D^{F}$, a similar approach to that of Section 3.2 can be used. If jointly valid marginal confidence intervals are constructed on each sample observation, calculating the estimated bounds, plugging lower confidence interval bounds into any lower bounds and upper confidence interval bounds into any upper bounds, will produce a confidence interval with correct coverage.
A.3 Sharp bounds given the joint distribution
To investigate how different the estimated bounds can be if we know the joint distribution of covariates $\mathbb{P}[X_{i}=x_{k}|G_{i}=g]$, as assumed in Cross and Manski (2002), I use the data sets from the three simulation exercises of Section 4.1 in Tables 1, 2, and 3. I calculated the lowest and highest possible lower and upper bounds under all different joint distributions for simulation exercises 1, 2, and 3 in Tables A.1, A.2, and A.3 respectively. In particular, I present the range of $L_{g}$ and $U_{g}$ over all $\{p_{kg}\}\in P_{g}$. The width of the range of possible lower and upper bounds tells us the extent to which knowing the joint distribution of covariates helps to narrow the identified set. In particular, the estimated bounds from my approach presented in Section 4.1 are exactly the lowest possible value of the lower bound and the highest possible value of the upper bound.
Table A.1: Estimated bounds under different joint covariate distributions, simulation exercise 1
Range of Lower Bounds Range of Upper Bounds
Parameter (1) (2)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.820,-0.580]$ $[0.531,0.860]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-0.997,-0.814]$ $[0.795,0.981]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-1,-0.946]$ $[0.947,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,-0.946]$ $[0.947,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$
Notes: See Section 4.1 for details about simulation exercise 1.
Table A.2: Estimated bounds under different joint covariate distributions, simulation exercise 2
Range of Lower Bounds Range of Upper Bounds
Parameter (1) (2)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.428,-0.369]$ $[0.688,0.749]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-1,-1]$ $[1,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-1,-0.988]$ $[0.988,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,-1]$ $[1,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$
Notes: See Section 4.1 for details about simulation exercise 2.
Table A.3: Estimated bounds under different joint covariate distributions, simulation exercise 3
Range of Lower Bounds Range of Upper Bounds
Parameter (1) (2)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.222,-0.225]$ $[0.768,0.860]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-1,-0.787]$ $[0.971,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-0.897,-0.775]$ $[0.972,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,-0.796]$ $[0.959,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$
Notes: See Section 4.1 for details about simulation exercise 3.
We see that knowing the joint distribution of the covariates helps make estimated bounds (weakly) narrower for all parameters. Bounds on certain parameters can become a lot narrower when we know the joint distribution, like the first parameter in simulation exercise 1. While knowing the joint distribution of the covariates will not help in signing any of the marginal effect parameters, these results suggests that bounds do look different when joint covariate information is unavailable in the aggregate data.
Appendix B Appendix: Proofs
B.1 Proof of Lemma 1
Proof.
As argued in the main text, the only information we have in addition to Assumption 1 are equations (1), (2), (3), and that $\delta_{k},\gamma_{kg},\pi_{kg}\in[0,1]$ for all $k,g$. The set given by (4) imposes all of these restrictions and nothing more, and finds the set of $\sum_{k=1}^{K}\lambda_{k}\delta_{k}$ such that the restrictions are satisfied. Thus the set is sharp. ∎
B.2 Proof of Proposition 1
Proof.
For any $g$, given any $\{p_{kg}\}_{k=1}^{K}$ such that $\sum_{k=1}^{K}p_{kg}=1$ and $p_{kg}\geq 0~{}\forall k$, note that there exists $\{c_{kg}\}_{k=1}^{K}\in[0,1]^{K}$ such that $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$; namely $c_{kg}=\mathbb{E}[Y_{i}|G_{i}=g]~{}\forall k$.
Thus imposing the restriction $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ does not further restrict the set of possible $\{p_{kg}\}_{k=1}^{K}$ if the following restrictions already hold: $\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}$, $\sum_{k=1}^{K}p_{kg}=1$ and $p_{kg}\geq 0~{}\forall k$. Finally note that all restrictions involving $\{p_{kg}\}_{k=1}^{K}$ only involve $p_{kg}$ with the same index $g$. This means it is equivalent to write (4) as
$D=\bigg{\{}\sum_{k=1}^{K}\lambda_{k}d_{k}\bigg{|}~{}0\leq d_{k}\leq 1~{}\forall k,\text{ and }\exists\,(p_{1g},\dots,p_{Kg})\in P_{g},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }d_{k}=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}~{}\forall k,~{}\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g\bigg{\}},$ (B.1)
where
$P_{g}=\bigg{\{}(p_{1g},\dots,p_{Kg})\in[0,1]^{K}\bigg{|}~{}\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}~{}\forall\ell,k,\text{ and }\sum_{k=1}^{K}p_{kg}=1\bigg{\}}.$ (B.2)
If $\mathbb{P}[G_{i}=g],c_{kg}\in[0,1]$ for all $k$ and $g$ then because $d_{k}=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}$ we know $d_{k}\in[0,1]$ for all $k$. Then we can get rid of $d_{k}$ in (B.1) by plugging in one of the constraints like so:
$D=\bigg{\{}\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}\bigg{|}~{}\exists\,(p_{1g},\dots,p_{Kg})\in P_{g},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g\bigg{\}}.$ (B.3)
I next show that $D$ is an interval. For any given $\{p_{kg}\}\in P_{g}$, we argued above that $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\mathbb{E}[Y_{i}|G_{i}=g]\in D$. For any $g$ consider arbitary $\{c_{kg}\},\{c_{kg}^{\prime}\}\in[0,1]^{K}$ with corresponding $\{p_{kg}\},\{p_{kg}^{\prime}\}\in P_{g}$ such that $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ and $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}^{\prime}p_{kg}^{\prime}$. We know $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}\in D$ and $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}^{\prime}\in D$. For arbitrary $t\in[0,1]$, let
$\tilde{c}_{kg}=tc_{kg}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$
for all $k,g$ and let
$\tilde{c}_{kg}^{\prime}=tc_{kg}^{\prime}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$
for all $k,g$. Then $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k}p_{kg}\tilde{c}_{kg}=\sum_{k}p_{kg}^{\prime}\tilde{c}_{kg}^{\prime}$ for all $g$ and $\tilde{c}_{kg},\tilde{c}_{kg}^{\prime}\in[0,1]$ for all $k,g$. Thus $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\tilde{c}_{kg}\in D$ and $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\tilde{c}_{kg}^{\prime}\in D$. Note
$\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\tilde{c}_{kg}$
$=t\left(\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}\right)+(1-t)\left(\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\mathbb{E}[Y_{i}|G_{i}=g]\right),$
$\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\tilde{c}_{kg}^{\prime}$
$=t\left(\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}^{\prime}\right)+(1-t)\left(\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\mathbb{E}[Y_{i}|G_{i}=g]\right).$
Since $t$ was arbitrary between 0 and 1, $D$ contains any value between $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\mathbb{E}[Y_{i}|G_{i}=g]$. Similarly, $D$ contains any value between $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}^{\prime}$ and $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\mathbb{E}[Y_{i}|G_{i}=g]$. In particular, $D$ contains any value between $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}^{\prime}$. Thus $D$ is an interval.
This means that we can equivalently express (B.3) and (B.2) as optimization problems as follows: $D=\left[L,U\right]$, where
$L$ $\equiv\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},$
$U$ $\equiv\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},\text{ and}$
$P_{g}=\underset{\{p_{kg}\}\in[0,1]^{K}}{\text{arg}\min}\sum_{r=1}^{LK}v_{r}^{+}+v_{r}^{-}\;\;\text{s.t. }\sum_{k=1}^{K}p_{kg}=1,~{}p_{kg},v_{r}^{+},v_{r}^{-}\geq 0~{}\forall k,r,\text{ and}\\[-10.0pt] \mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]-\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}=v_{K(\ell-1)+k}^{+}-v_{K(\ell-1)+k}^{-}~{}\forall\ell,k.$
The slack variables in the $P_{g}$ problem will all be equal to zero at optimum.
Finally, since the constraints in $L$ and $U$ are for each $g$, we can equivalently solve a program for each group $g$ and take a weighted sum of the solutions to get lower and upper bounds for $D$, like so: $D=\left[\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]L_{g},\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]U_{g}\right]$, where
$L_{g}$ $\equiv\min_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},$
$U_{g}$ $\equiv\max_{\{c_{kg}\}\in[0,1]^{K}}\sum_{k=1}^{K}\lambda_{k}c_{kg}\text{ s.t. }\exists\{p_{kg}\}\in P_{g}\text{ with }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}.$
B.3 Proof of Proposition 2
Proof.
To prove sharpness of the identified set, it is sufficient to show that the following set is an interval: for $c_{g}\equiv(c_{1g},\dots,c_{Kg})^{\prime}$,
$\bigg{\{}\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}\bigg{|}~{}S_{g}c_{g}\leq a_{g}~{}\forall g\text{ and }\exists\,(p_{1g},\dots,p_{Kg})\in P_{g},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g\bigg{\}}.$ (B.4)
We know that the set of $c_{g}\in[0,1]^{KG}$ such that $S_{g}c_{g}\leq a_{g}$ for all $g$ is a convex, closed, and bounded set because it is the intersection of a closed, bounded, and convex set with a half-space. This means the set of values of $\sum_{k}\lambda_{k}c_{kg}$ such that $S_{g}c_{g}\leq a_{g}$ for all $g$ is also a closed and bounded and convex set.
In the proof of Proposition 1 I argued that for any $c_{g},c_{g}^{\prime}\in[0,1]^{K}$ with corresponding $\{p_{kg}\},\{p_{kg}^{\prime}\}\in P_{g}$ that satisfy
$\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ $\qquad\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}^{\prime}p_{kg}^{\prime}$ (B.5)
$S_{g}c_{g}\leq a_{g}$ $\qquad S_{g}c_{g}^{\prime}\leq a_{g},$ (B.6)
we know $tc_{kg}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$ and $tc_{kg}^{\prime}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$ also satisfy (B.5) in place of $c_{kg}$ and $c_{kg}^{\prime}$ respectively, for any $t\in[0,1]$. Thus for any $\sum_{k=1}^{K}\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}c_{kg}^{\prime}$, $\{\tilde{c}_{kg}\}$ satisfies $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}\tilde{c}_{kg}p_{kg}$. And that (B.6) holds implies $S_{g}(tc_{g}+(1-t)c_{g}^{\prime})\leq a_{g}$ for any $t\in[0,1]$, meaning that for any $\sum_{k=1}^{K}\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}c_{kg}^{\prime}$, $\{\tilde{c}_{kg}\}$ satisfies $S_{g}\tilde{c}_{g}\leq a_{g}$.
Thus any $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}^{\prime}$ is in the set given by (B.4), meaning (B.4) is an interval. Thus the sharp identified set is an interval. The rest of Proposition 2 can be proved following arguments similar to those used in the proof of Proposition 1. ∎
B.4 Proof of Proposition 3
Proof.
To prove sharpness of the identified set, it is sufficient to show that the following set is an interval:
$\bigg{\{}\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}\bigg{|}~{}\exists\,\{p_{kg}\}\in P_{g},\{c_{kg}\}\in[0,1]^{K}~{}\forall g\text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g,\\ \mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g},\forall g\bigg{\}}.$ (B.7)
Note that the set of $\{c_{kg}\}\in[0,1]^{KG}$ such that
$\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g},\forall g$ (B.8)
is a convex, closed, and bounded set because it is the intersection of a closed, bounded, and convex half-space. This means the set of values of $\sum_{k}\lambda_{k}c_{kg}$ such that (B.8) holds is also a closed and bounded and convex set.
In the proof of Proposition 1 I argued that for any $\{c_{kg}\},\{c_{kg}^{\prime}\}\in[0,1]^{K}$ with corresponding $\{p_{kg}\},\{p_{kg}^{\prime}\}\in P_{g}$ that satisfy
$\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},\qquad\qquad$ $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}^{\prime}p_{kg}^{\prime},$ (B.9)
$\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ $=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg},$ (B.10)
$\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ $=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}^{\prime}p_{kg}^{\prime},$ (B.11)
we know $tc_{kg}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$ and $tc_{kg}^{\prime}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$ also satisfy (B.9) in place of $c_{kg}$ and $c_{kg}^{\prime}$ respectively, for any $t\in[0,1]$. Thus for any $\sum_{k=1}^{K}\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}c_{kg}^{\prime}$, $\{\tilde{c}_{kg}\}$ satisfies $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}\tilde{c}_{kg}p_{kg}$.
Note that, similar to the claim made in the proof of Proposition 1, since $\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{kg}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{kg}^{\prime}$ it follows that (B.10) still holds when letting $c_{kg}=\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$ and (B.11) still holds when letting $c_{kg}^{\prime}=\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$.
Thus $tc_{kg}+(1-t)\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$ in place of $c_{kg}$ and $tc_{kg}^{\prime}+(1-t)\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$ in place of $c_{kg}^{\prime}$ also satisfy (B.10) and (B.11) respectively for any $t\in[0,1]$. So for any $\sum_{k=1}^{K}\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}c_{kg}^{\prime}$, $\{\tilde{c}_{kg}\}$ in place of $\{c_{kg}\}$ satisfies (B.10).
Thus any $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}^{\prime}$ is in the set given by (B.7), meaning (B.7) is an interval. Thus the sharp identified set is an interval. The rest of Proposition 3 can be proved following arguments similar to those used in the proof of Proposition 1. ∎
B.5 Proof of Proposition 4
Proof.
I prove the result for $\hat{D}$ defined with respect to Proposition 1, then discuss the additional restrictions implied by Assumptions 2 and 3.
The result follows from the Theorem of the Maximum for the following correspondence:
$\Gamma\left(y,\{g_{k,\ell}\}_{k,\ell}\right)=\underset{\{p_{k},c_{k}\}_{k}}{\text{arg}\max}\left(\underset{\{p_{k},c_{k}\}_{k}}{\text{arg}\min}\right)\bigg{\{}\sum_{k=1}^{K}\lambda_{k}c_{k}\text{ s.t. }y=\sum_{k=1}^{K}c_{k}p_{k},\sum_{k=1}^{K}p_{k}=1,\\ \qquad g_{k,\ell}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j},c_{k}\in[0,1]~{}\forall k,p_{k}\geq 0~{}\forall k\bigg{\}}.$
Since the $\widehat{Pr}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ are valid marginal distributions for $X_{i}$ with respect to the assumed support, this correspondence is analogous to the set defined in Proposition 2; that is, the constraints in $P_{g}$ hold with $v_{r}^{+}=v_{r}^{-}=0$.
To apply the Theorem of the Maximum, I must show $\Gamma$ is a continuous and compact-valued correspondence. That $\Gamma$ is compact-valued follows because the constraints are intersections of non-parallel planes with a curve $(y=\sum_{k=1}^{K}c_{k}p_{k})$ on a compact set $(c_{k}\in[0,1],p_{k}\geq 0,\sum_{k=1}^{K}p_{k}=1)$.
Let $(y^{n},\{g_{k,\ell}^{n}\}_{k,\ell})$ be an arbitrary sequence such that $(y^{n},\{g_{k,\ell}^{n}\}_{\ell,j})\to(y,\{g_{k,\ell}\}_{k,\ell})$ as $n\to\infty$ and all $y^{n},y\in[0,1]$ and $\{g_{k,\ell}^{n}\}_{k,\ell}$ are valid marginal distributions given the assumed support. Let ($(\{p_{k}^{n}\},\{c_{k}^{n}\})$ be an arbitrary sequence such that ($(\{p_{k}^{n}\},\{c_{k}^{n}\})\to(\{p_{k}\},\{c_{k}\})$ as $n\to\infty$ and for each $n$, ($(\{p_{k}^{n}\},\{c_{k}^{n}\})\in\Gamma\left((y^{n},\{g_{k,\ell}^{n}\}_{k,\ell})\right)$. If we can show ($(\{p_{k}\},\{c_{k}\})\in\Gamma((y,\{g_{k,\ell}\}_{k,\ell}))$ then $\Gamma$ is upper hemicontinuous.
Since $\sum_{k=1}^{K}p_{k}^{n}=1$ for all $n$, $p_{k}^{n}\geq 0$ for all $k,n$, and $p_{k}^{n}\to p_{k}$ for all $k$, it follows that $\sum_{k=1}^{K}p_{k}=1$ and $p_{k}\geq 0$ for all $k$. Since $c_{k}^{n}\in[0,1]$ for all $k,n$ and $c_{k}^{n}\to c_{k}$ for all $k$, it follows that $c_{k}\in[0,1]$ for all $k$. Since $y^{n}=\sum_{k=1}^{K}p_{k}^{n}c_{k}^{n}$ for all $k,n$ and $y^{n}\to y,d_{k}^{n}\to d_{k},p_{k}^{n}\to p_{k}$ for all $k,n$, it follows that $y=\sum_{k}p_{k}d_{k}$. Since $g_{k,\ell}^{n}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}^{n}$ for all $k,\ell,n$ and we have $p_{k}^{n}\to p_{k},g_{k,\ell}^{n}\to g_{k,\ell}$ for all $k,\ell$, it follows that $g_{k,\ell}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}$ for all $k,\ell$. Thus ($(\{p_{k}\},\{c_{k}\})\in\Gamma((y,\{g_{k,\ell}\}_{k,\ell}))$ and thus $\Gamma$ is upper hemicontinuous.
Let $(y^{n},\{g_{k,\ell}^{n}\}_{k,\ell})$ be an arbitrary sequence such that $(y^{n},\{g_{k,\ell}^{n}\}_{\ell,j})\to(y,\{g_{k,\ell}\}_{k,\ell})$ as $n\to\infty$ and all $y^{n},y\in[0,1]$ and $\{g_{k,\ell}^{n}\}_{k,\ell}$ are valid marginal distributions given the assumed support. Let ($(\{p_{k}\},\{c_{k}\})\in\Gamma(y,\{g_{k,\ell}\}_{k,\ell})$ be arbitrary. If we can show there exists a subsequence $(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$ and sequence ($(\{p_{k}^{t}\},\{c_{k}^{t}\})$ such that ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\to(\{p_{k}\},\{c_{k}\})$ as $t\to\infty$ and ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\in\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$ for all $t$, then $\Gamma$ is lower hemicontinuous.
We know that $g_{k,\ell}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}$ and $\sum_{k=1}^{K}p_{k}=1$. Because we assume each $\{g_{k,\ell}\}$ is a valid marginal distribution of covariates there are strictly less than $K$ unique noncollinear equations in this system of equations. We can express these equations as a linear system $Ap=g$, where $p=\{p_{k}\}$ and $g=\{g_{k,\ell}\}_{k,\ell}$, and all solutions are given by $p=A^{+}g+(I-A^{+}A)w$, where $A^{+}$ is the Moore-Penrose inverse and $w$ is any arbitrary vector of correct dimension.
Thus the solution $\{p_{k}\}$ to this system of linear equations is continuous in $\{g_{k,\ell}\}_{k,\ell}$. Therefore for each $t$ there exists $n_{t}$ large enough that we can find $\{p_{k}^{t}\}$ arbitrarily close to $\{p_{k}\}$ such that $g_{k,\ell}^{n_{t}}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}^{t}$ and $\sum_{k=1}^{K}p_{k}^{t}=1$. Such a solution exists because we assume $\hat{D}$ is always nonempty with a valid marginal distribution of covariates $\{g_{k,\ell}^{n_{t}}\}$.
Recall each $p_{k}\geq 0$. In fact, the proposition imposes as an assumption that we only need consider $\{p_{k}\}$ such that each $p_{k}>0$. Thus we can make $p_{k}^{t}\geq 0$ for each $k$. Thus we can construct a sequence $\{p_{k}^{t}\}$ with corresponding subsequence $\{g_{k,\ell}^{n_{t}}\}_{k,\ell}$ such that $g_{k,\ell}^{n_{t}}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}^{t}$ for all $t$, $\sum_{k=1}^{K}p_{k}^{t}=1$ for all $t$, $p_{k}^{t}\geq 0$ for all $k$, and $p_{k}^{t}\to p_{k}$ for all $k$ as $t\to\infty$.
Given this sequence $\{p_{k}^{t}\}$, we need to find a corresponding sequence $\{c_{k}^{t}\}$ and subsequence $\{y^{n_{t}}\}$ such that $y^{n_{t}}=\sum_{k=1}^{K}c_{k}^{t}p_{k}^{t}$ and $c_{k}^{t}\in[0,1]$ for all $k,t$. Note that we can write the equation $y=\sum_{k=1}^{K}c_{k}p_{k}$ as $p^{\prime}c=y$ and consider the Moore-Penrose representation of the set of all solutions $c$. As a matrix algebra result, it is true that the Moore-Penrose inverse of $p$ is continuous, that is $(p_{t})^{+}\to p^{+}$ as long as the rank of $p_{t}$ is the same as the rank of $p$ for all $t$. Since the rank of $p_{t}$ and $p$ is always 1, it follows that the solution $c$ to the equation is continuous in $p$ and $y$.
Therefore for each $t$ there exists $n_{t}$ large enough that we can find $\{\tilde{c}_{k}^{t}\}$ arbitrarily close to $\{c_{k}\}$ with $y^{n_{t}}=\sum_{k=1}^{K}\tilde{c}_{k}^{t}p_{k}^{t}$ (with some relabeling of the $\{p_{k}^{t}\}$ sequence indices as necessary). We know $c_{k}\in[0,1]$ for each $k$, but it may be the case that $\tilde{c}_{k}^{t}\notin[0,1]$ for some $k$. However for those $k$ we will have $\tilde{c}_{k}^{t}$ arbitrarily close to $[0,1]$. As argued in the proof of Proposition 1 it also holds that $y^{n_{t}}=\sum_{k=1}^{K}y^{n_{t}}p_{k}^{t}$ (and we know $y^{n,t}\in[0,1]$), meaning there exists another feasible $\{c_{k}^{t}\}$ between $\{\tilde{c}_{k}^{t}\}$ and $(y^{n_{t}},\dots,y^{n_{t}})$ that is arbitrarily close to $\{\tilde{c}_{k}^{t}\}$ but with $c_{k}^{t}\in[0,1]$ for all $k$.
Thus we have sequence ($(\{p_{k}^{t}\},\{c_{k}^{t}\})$ with corresponding subsequence $(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$ such that ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\to(\{p_{k}\},\{c_{k}\})$ as $t\to\infty$, and ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\in\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$. Therefore $\Gamma$ is lower hemicontinuous.
[…]
The law of large numbers also gives us that $\widehat{Pr}[G_{i}=g]$ is consistent, so that by continuous mapping theorem again we have that the lower and upper bounds of $\hat{D}$ converge to the lower and upper bounds of $D$.
To accommodate the restriction that $S_{g}c\leq a_{g}$ under Assumption 2 in correspondence $\Gamma$, note that any sequence $\{c_{k}^{n}\}\to\{c_{k}\}$ with ($(\{p_{k}^{n}\},\{c_{k}^{n}\})\in\Gamma\left((y^{n},\{g_{k,\ell}^{n}\}_{k,\ell})\right)$ satisfies $S_{g}c^{n}\leq a_{g}$ and thus $S_{g}c\leq a_{g}$ by continuity as well. So upper hemicontinuity is maintained.
The assumption that $\hat{D}$ is nonempty and the observed marginals are valid with respect to the assumed support ensures that for $(y^{n},\{g^{n}_{k,\ell}\}_{k,\ell})\to(y,\{g_{k,\ell}\}_{k,\ell})$ there exists ($(\{p_{k}\},\{c_{k}\})\in\Gamma(y,\{g_{k,\ell}\}_{k,\ell})$ and that $\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$ is nonempty. Let ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\in\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$ be a sequence such that $\max_{k}\{|p_{k}^{t}-p_{k}|,|c_{k}^{t}-c_{k}|\}$ is minimized over ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\in\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$ for each $t$.
I claim ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\to(\{p_{k}\},\{c_{k}\})$. From the restrictions that $g_{k,\ell}^{n_{t}}=\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}^{t}$ we know $\max_{k}\sum_{k=1}^{K}|p_{k}^{t}-p_{k}|\to 0$ as discussed above with the Moore-Penrose inverse solution to the linear system, so $\{p_{k}^{t}\}\to\{p_{k}\}$. And if $\max_{k}\sum_{k=1}^{K}|c_{k}^{t}-c_{k}|\not\to 0$ then $c_{k}^{t}p_{k}^{t}\not\to c_{k}p_{k}$, meaning the constraint $y=\sum_{k=1}^{K}c_{k}p_{k}$ cannot hold either as $y^{n_{t}}\to y$. Thus it must be that $\{c_{k}^{t}\}\to\{c_{k}\}$ and so lower hemicontinuity is also maintained because $S_{g}c^{t}\leq a_{g}$ by assumption.
To accommodate the restriction
$y_{k,\ell}g_{k,\ell}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{j}p_{j}$ (B.12)
under Assumption 3 in correspondence $\Gamma$, note that any sequence ($(\{p_{k}^{n}\},\{c_{k}^{n}\})\to(\{p_{k}\},\{c_{k}\})$ with ($(\{p_{k}^{n}\},\{c_{k}^{n}\})\in\Gamma\left((y^{n},\{g_{k,\ell}^{n}\}_{k,\ell},\{y_{k,\ell}^{n}\}_{k,\ell})\right)$ satisfies
$y_{k,\ell}^{n}g_{k,\ell}^{n}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{j}^{n}p_{j}^{n},$
and thus (B.12) holds by continuity as well. So upper hemicontinuity is maintained.
The assumption that $\hat{D}$ is nonempty and the observed marginals are valid with respect to the assumed support ensures that for $(y^{n},\{g^{n}_{k,\ell}\}_{k,\ell},\{y^{n}_{k,\ell}\}_{k,\ell})\to(y,\{g_{k,\ell}\}_{k,\ell},\{y_{k,\ell}\}_{k,\ell})$ there exists ($(\{p_{k}\},\{c_{k}\})\in\Gamma(y,\{g_{k,\ell}\}_{k,\ell},\{y_{k,\ell}\}_{k,\ell})$ and that $\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell},\{y_{k,\ell}^{n_{t}}\}_{k,\ell})$ is nonempty. Let ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\in\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell},\{y_{k,\ell}^{n_{t}}\}_{k,\ell})$ be a sequence such that $\max_{k}\{|p_{k}^{t}-p_{k}|,|c_{k}^{t}-c_{k}|\}$ is minimized over ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\in\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$for each $t$.
I claim ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\to(\{p_{k}\},\{c_{k}\})$. From the restrictions that $g_{k,\ell}^{n_{t}}=\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}^{t}$ we know $\max_{k}\sum_{k=1}^{K}|p_{k}^{t}-p_{k}|\to 0$ as discussed above with the Moore-Penrose inverse solution to the linear system. And if $\max_{k}\sum_{k=1}^{K}|c_{k}^{t}-c_{k}|\not\to 0$ then $c_{k}^{t}p_{k}^{t}\not\to c_{k}p_{k}$, meaning the constraints $y=\sum_{k=1}^{K}c_{k}p_{k}$ and $y_{k,\ell}g_{k,\ell}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{j}p_{j}$ cannot hold either as $y^{n_{t}}\to y,g_{k,\ell}^{n_{t}}\to g_{k,\ell},y_{k,\ell}^{n_{t}}\to y_{k,\ell}$. Thus it must be that $\{c_{k}^{t}\}\to\{c_{k}\}$ and so lower hemicontinuity is also maintained. ∎
B.6 Proof of Proposition 5
Proof.
Since Clopper-Pearson intervals are finite-sample valid for each sample observation, using the Bonferroni correction means that the population analog of every single sample observation is contained in its Clopper-Pearson interval with joint probability greater than $1-\alpha$.
Note that if all population analogs of the sample observations are jointly in their respective Clopper-Pearson intervals then any $\{c_{kg}\}$ and $\{p_{kg}\}$ consistent with the population $P_{g}$ and $L_{g},U_{g}$ constraints in the formulation of $D$ are also consistent with the $\hat{P}_{g,CI}$ and $\hat{L}_{g,CI},\hat{U}_{g,CI}$ constraints in the formulation of $\hat{D}_{CI}$. Thus the set of $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ consistent with the population $P_{g}$ and $L_{g},U_{g}$ constraints in the formulation of $D$ is a subset of the set of $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ consistent with the $\hat{P}_{g,CI}$ and $\hat{L}_{g,CI},\hat{U}_{g,CI}$ constraints in the formulation of $\hat{D}_{CI}$. And if the true $\mathbb{P}[G_{i}=g]$ are all contained in their Clopper-Pearson intervals, it follows that $D\subseteq\hat{D}_{CI}$.
Since the event that all population analogs of the sample observations are jointly in their respective Clopper-Pearson intervals happens with probability at least $1-\alpha$, it follows that $\mathbb{P}[D\subseteq\hat{D}_{CI}]\geq 1-\alpha$. ∎
B.7 Proof of Corollary A.1
Proof.
For given weights $\lambda_{1},\dots,\lambda_{K}$ and any fixed $\{p_{kg}\}\in P_{g}$, relabel the indices $k=1,\dots,K$ so that $\frac{\lambda_{1}}{p_{1g}}\geq\dots\geq\frac{\lambda_{K}}{p_{Kg}}$, where if $p_{kg}=0$ we define $\frac{\lambda_{k}}{p_{gk}}\equiv+\infty$.
Note that the proposed solution in Corollary A.1 is feasible. We will take advantage of strong duality (because Slater’s condition holds) and use the joint feasiblity and satisfaction of complementary slackness for proposed solutions to the primal and dual problems to show optimality.
The dual of the minimization linear program is
$\max_{u\in\mathbb{R},v\in\mathbb{R}^{2K}}-\mathbb{E}[Y_{i}|G_{i}=g]u-\sum_{i=1}^{K}v_{i}\text{ s.t. }-p_{kg}u-v_{k}\leq\lambda_{k}~{}\forall k=1,\dots,K,v_{i}\geq 0~{}\forall i=1,\dots,K.$
Let $k$ be such that $\sum_{j=1}^{k-1}p_{jg}\leq 1-\mathbb{E}[Y_{i}|G_{i}=g]<\sum_{j=1}^{k}p_{jg}$ holds. Consider a solution to the dual where $v_{i}=0$ for all $i\leq k$, and for $i\geq k$ we have that $v_{i}$ satisfies $-p_{ig}u-v_{i}=\lambda_{i}$, meaning $u=-\frac{\lambda_{k}}{p_{kg}}$. Thus for $i>k$ we have $v_{i}>0$ because $-p_{ig}u=\frac{\lambda_{k}p_{ig}}{p_{kg}}\geq\lambda_{i}$. Clearly this is a feasible solution.
We see that complementary slackness holds with the condition that $c_{ig}\leq 1$ for all $i$ and $v_{i}\geq 0$ for all $i$ because $c_{ig}\neq 1$ for all $i\leq k$ while $v_{i}=0$ for all $i\leq k$. Complementary slackness holds with the condition that $c_{ig}\geq 0$ for all $i$ and $-p_{ig}u-v_{i}\leq\lambda_{i}$ for all $i$ because $c_{ig}\neq 0$ for all $i\geq k$ while $-p_{ig}u-v_{i}=\lambda_{i}$ for all $i\geq k$.
Thus the proposed solution for the minimization problem is optimal.
The dual of the maximization linear program is
$\min_{u\in\mathbb{R},v\in\mathbb{R}^{2K}}\mathbb{E}[Y_{i}|G_{i}=g]u+\sum_{i=1}^{K}v_{i}\text{ s.t. }p_{k}u+v_{k}\geq\lambda_{k}~{}\forall k=1,\dots,K,v_{i}\geq 0~{}\forall i=1,\dots,K.$
Let $k$ be such that $\sum_{j=1}^{k-1}p_{jg}\leq\mathbb{E}[Y_{i}|G_{i}=g]<\sum_{j=1}^{k}p_{jg}$ holds. Consider a solution to the dual where $v_{i}=0$ for all $i\geq k$, and for $i\leq k$ we have that $v_{i}$ satisfies $p_{ig}u+v_{i}=\lambda_{i}$, meaning $u=\frac{\lambda_{k}}{p_{kg}}$. Thus for $i<k$ we have $v_{i}>0$ because $p_{ig}u=\frac{\lambda_{k}p_{ig}}{p_{kg}}\leq\lambda_{i}$. Clearly this is a feasible solution.
We see that complementary slackness holds with the condition that $c_{ig}\leq 1$ for all $i$ and $v_{i}\geq 0$ for all $i$ because $c_{ig}\neq 1$ for all $i\geq k$ while $v_{i}=0$ for all $i\geq k$. Complementary slackness holds with the condition that $c_{ig}\geq 0$ for all $i$ and $p_{ig}u+v_{i}\geq\lambda_{i}$ for all $i$ because $c_{ig}\neq 0$ for all $i\leq k$ while $p_{ig}u+v_{i}=\lambda_{i}$ for all $i\leq k$.
Thus the proposed solution for the minimization problem is optimal. ∎
B.8 Proof of Proposition A.1
Proof.
First I show that $D\subseteq D^{F}$.
Since $D$ is truly an interval, as proved when proving Proposition 1, there exists some $\{c_{kg}\}_{k,g},\{p_{kg}\}_{k,g}$ that satisfy the constraints of $L_{g},U_{g},$ and $P_{g}$ where $\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\sum_{k=1}^{K}\lambda_{k}c_{kg}\in D$.
By an argument analogous to that given in the proof of Proposition 1, the set of each $c_{kg}$ that satisfy the constraints is the sharp identified set for $\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]$. Thus it is sufficient to show that for each $k,g$,
$\frac{L_{g}(1,x_{k})}{L_{g}(1,x_{k})+U_{g}(0,x_{k})}\leq\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]\leq\frac{U_{g}(1,x_{k})}{U_{g}(1,x_{k})+L_{g}(0,x_{k})}.$
This means $\left[\frac{L_{g}(1,x_{k})}{L_{g}(1,x_{k})+U_{g}(0,x_{k})},\frac{U_{g}(1,x_{k})}{U_{g}(1,x_{k})+L_{g}(0,x_{k})}\right]$ are also bounds, so it follows that
$\frac{L_{g}(1,x_{k})}{L_{g}(1,x_{k})+U_{g}(0,x_{k})}\leq c_{kg}\leq\frac{U_{g}(1,x_{k})}{U_{g}(1,x_{k})+L_{g}(0,x_{k})}.$
and thus $\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\sum_{k=1}^{K}\lambda_{k}c_{kg}\in D^{F}$.
Note
$\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]$ $=\mathbb{P}[Y_{i}=1|X_{i}=x_{k},G_{i}=g]$
$=\frac{\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]}{\mathbb{P}[X_{i}=x_{k}|G_{i}=g]}$
$=\frac{\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]}{\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]+\mathbb{P}[Y_{i}=0,X_{i}=x_{k}|G_{i}=g]}.$
One can check that the function $\frac{x}{x+y}$ is increasing in $x$ and decreasing in $y$; thus $\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]$ attains its minimum when $\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]$ is as small as possible and $\mathbb{P}[Y_{i}=0,X_{i}=x_{k}|G_{i}=g]$ is as large as possible. Similarly $\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]$ attains its maximum when $\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]$ is as large as possible and $\mathbb{P}[Y_{i}=0,X_{i}=x_{k}|G_{i}=g]$ is as small as possible.
Since $\mathbb{P}[Y_{i}=0,X_{i}=x_{k}|G_{i}=g]\in[L_{g}(0,x_{k}),U_{g}(0,x_{k})]$ and $\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]\in[L_{g}(1,x_{k}),U_{g}(1,x_{k})]$, it follows then that $\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]\in\left[\frac{L_{g}(1,x_{k})}{L_{g}(1,x_{k})+U_{g}(0,x_{k})},\frac{U_{g}(1,x_{k})}{U_{g}(1,x_{k})+L_{g}(0,x_{k})}\right]$.
Now note that if the Fréchet inequalities on $\mathbb{P}[Y_{i}=y,X_{i}=x_{k}|G_{i}=g]$ are sharp for all $y\in\{0,1\},k=1,\dots,K,g=1,\dots,G$ then $\mathbb{P}[Y_{i}=0,X_{i}=x_{k}|G_{i}=g]$ attains its minimum at $L_{g}(0,x_{k})$ and its maximum at $U_{g}(0,x_{k})$, and $\mathbb{P}[Y_{i}=1,X_{i}=x_{k}|G_{i}=g]$ attains its minimum at $L_{g}(1,x_{k})$ and its maximum at $U_{g}(1,x_{k})$. Thus bounds $\left[\frac{L_{g}(1,x_{k})}{L_{g}(1,x_{k})+U_{g}(0,x_{k})},\frac{U_{g}(1,x_{k})}{U_{g}(1,x_{k})+L_{g}(0,x_{k})}\right]$ on $\mathbb{E}[Y_{i}|X_{i}=x_{k},G_{i}=g]$ are indeed sharp and so $D=D^{F}$. ∎